### v1

# ConDiSim: Conditional Diffusion Models for Simulation-Based Inference

Abstract
We present a conditional diffusion model - ConDiSim, for simulation-based inference of complex systems with intractable likelihoods. ConDiSim leverages denoising diffusion probabilistic models to approximate posterior distributions, consisting of a forward process that adds Gaussian noise to parameters, and a reverse process learning to denoise, conditioned on observed data. This approach effectively captures complex dependencies and multi-modalities within posteriors. ConDiSim is evaluated across ten benchmark problems and two real-world test problems, where it demonstrates effective posterior approximation accuracy while maintaining computational efficiency and stability in model training. ConDiSim offers a robust and extensible framework for simulation-based inference, particularly suitable for parameter inference workflows requiring fast inference methods.
1 Introduction
Statistical inference of model parameters from empirical observations is a fundamental challenge in scientific research, enabling researchers to derive meaningful insights from complex simulation models. These parameters govern the behavior of simulators that replicate real-world phenomena, providing a bridge between theoretical constructs and empirical observations [Lavin et al., 2021]. Calibrating these parameters to ensure that simulator outputs align with observed data constitutes an inverse problem, formally defined within the framework of simulation-based inference (SBI) [Cranmer et al., 2020]. Solving this inverse problem involves addressing uncertainties arising from model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations or similar parameters may lead to varied outputs. Additionally, parameter inference becomes increasingly complex when simulators operate as ‘black boxes’ with intractable likelihood functions, rendering traditional likelihood-based Bayesian methods impractical [Sisson et al., 2018].
Recent advances in machine learning have introduced powerful generative models capable of approximating complex data distributions without explicit likelihood computations. Among these, diffusion models have emerged as a particularly promising class [Sohl-Dickstein et al., 2015]. Originally developed for image and signal processing, diffusion models iteratively denoise data starting from pure noise, effectively learning data distributions through a sequence of transformations. Compared to alternative generative frameworks like Variational Autoencoders (VAEs) [Kingma and Welling, 2014], Generative Adversarial Networks (GANs) [Goodfellow et al., 2014], and flow-based models [Rezende and Mohamed, 2015], diffusion models offer significant advantages. They avoid issues like posterior collapse in VAEs and mode collapse in GANs, and are not constrained by invertible transformations as in flow-based models, enabling them to better model complex, high-dimensional, and multi-modal distributions. Diffusion models also demonstrate superior training stability and come with theoretical convergence guarantees due to their foundation in stochastic differential equations [Ho et al., 2020]. Their flexible architectures and sampling processes allow practitioners to balance sample quality and generation speed, enhancing their applicability in various contexts, including SBI.
Applying diffusion models to SBI leverages their ability to implicitly learn the conditional distribution of simulator parameters given observations, circumventing the need for explicit likelihood evaluations. This is particularly advantageous for inverse problems involving black-box simulators. Moreover, diffusion models naturally support amortized inference: a single, globally trained model approximates the posterior distribution across multiple observations, eliminating the need for separate optimization for each new observation [Kingma and Welling, 2014]. This capability is especially valuable when rapid inference is required. However, integrating diffusion models into SBI presents challenges. Their iterative sampling process can be computationally intensive, and their performance is sensitive to hyperparameters like noise schedules and network architecture. Addressing these issues is crucial for enhancing the practical utility of diffusion-based SBI methods [Nichol and Dhariwal, 2021].
Contribution: We present a conditional diffusion model (ConDiSim) for simulation-based inference that balances model expressiveness with training efficiency and simplicity. Experiments on benchmark tasks and real-world datasets show the effectiveness of the approach in recovering complex posterior distributions, while incurring minimal computational overhead. Further, simulation-based calibration checks validate the reliability and robustness of the proposed parameter inference approach.
The paper is organized as follows. Section 2 outlines the related work. In Section 3, we formally define the SBI problem, followed by a detailed explanation of the ConDiSim model in Section 3.1 and Section 3.2 details our amortized inference strategy. A comprehensive comparison on benchmark, and real-world test problems is presented in Section 4, with additional results provided in the Appendix. Section 5 discusses the results, while Section 6 summarizes the key findings and suggests future directions.
2 Related Work
Simulation-based inference (SBI) has advanced significantly in recent years, evolving from traditional likelihood-free methods to modern deep learning approaches. Classical techniques like Approximate Bayesian Computation (ABC) [Beaumont et al., 2002] and its sequential variants such as Sequential Monte Carlo-ABC (SMC-ABC) [Toni et al., 2009], provide mathematical convergence guarantees but often struggle with high-dimensional problems [Sisson et al., 2018]. Although neural network-learned summary statistics [Jiang et al., 2017, Wrede et al., 2022] have contributed to mitigating this limitation, they introduce additional computational complexity.
Early advancements in neural network-based amortized inference for SBI were pioneered by Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016], which introduced conditional normalizing flows to model the posterior distribution, significantly improving accuracy and scalability over traditional methods. Sequential extensions like SNPE-A [Papamakarios and Murray, 2018], SNPE-B [Lueckmann et al., 2017], and SNPE-C (also known as Automatic Posterior Transformation or APT) [Greenberg et al., 2019] enhanced efficiency by iteratively refining the model to focus on relevant parameter regions. However, these sequential methods require multiple training rounds and are constrained by the invertible transformations required in flow-based models.
Alternative sequential algorithms have also been proposed, such as Sequential Neural Ratio Estimation (SNRE) [Hermans et al., 2020], which learns the likelihood-to-evidence ratio, and Sequential Neural Likelihood Estimation (SNLE) [Papamakarios et al., 2019], which approximates the likelihood function directly. Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] leverages adversarial networks to model complex distributions without explicit likelihood calculations. While effective for challenging inference problems, GATSBI faces issues common to adversarial training, such as instability and mode collapse [Arjovsky and Bottou, 2017, Arjovsky et al., 2017], which can limit robustness and complicate training, and it struggles in low-dimensional spaces as noted in [Ramesh et al., 2022].
Recently, Simformer [Gloeckler et al., 2024] combined score-based diffusion models with transformer architectures to accurately learn complex, high-dimensional posteriors, offering high flexibility albeit with substantial computational overhead. In contrast, our proposed method, ConDiSim, utilizes denoising-based diffusion models with a simpler architecture, maintaining competitive performance while being computationally efficient. This makes it suitable for applications such as model exploration, that may involve numerous parameter inference problems with inference time being a key consideration [Wrede and Hellander, 2019].
In this study, we evaluate CondiSim against four distinct simulation-based inference (SBI) approaches: GATSBI [Ramesh et al., 2022], APT [Greenberg et al., 2019], NPE [Papamakarios and Murray, 2016], and Simformer [Gloeckler et al., 2024]. These methods represent a diverse spectrum of generative modeling techniques—namely, flow-based, adversarial, diffusion, one-shot, and sequential frameworks—adapted for SBI. This comprehensive selection facilitates an evaluation that emphasizes efficiency and scalability rather than focusing solely on the accuracy of posterior approximations. For further details on deep learning methodologies applied to SBI, readers are referred to [Zammit-Mangion et al., 2024, Radev et al., 2023a].
3 Simulation-Based Inference
Simulation-based inference (SBI) aims to infer the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ of model parameters $\bm{\theta}$, given observed data $\mathbf{y}$. From Bayes’ theorem:
$p(\bm{\theta}\mid\mathbf{y})\propto p(\mathbf{y}\mid\bm{\theta})\,p(\bm{\theta}),$ (1)
where $p(\mathbf{y}\mid\bm{\theta})$ is the likelihood function and $p(\bm{\theta})$ is the prior distribution. In many complex models, especially those involving stochastic simulators, the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is intractable or computationally expensive to evaluate, rendering traditional inference methods like Markov Chain Monte Carlo (MCMC) algorithms [Andrieu et al., 2003] infeasible due to their reliance on explicit likelihood evaluations.
When the likelihood function is inaccessible, a popular strategy is to generate empirical samples using a stochastic simulator $\mathcal{M}(\bm{\theta},\bm{\xi})$ that produces synthetic data $\mathbf{y}$ given parameters $\bm{\theta}$ and random variables $\bm{\xi}$ representing system noise. Sampling $\bm{\theta}$ from a prior distribution $p(\bm{\theta})$ informed by domain knowledge or chosen to be noninformative and generating the corresponding $\mathbf{y}$ through the simulator, we obtain samples from the joint distribution $p(\bm{\theta},\mathbf{y})$. These joint samples allow us to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ using generative modeling techniques that learn the inverse mapping from observations $\mathbf{y}$ to parameters $\bm{\theta}$ without requiring explicit likelihood calculations.
3.1 ConDiSim: A Conditional Diffusion Framework for Simulation-Based Inference
We now introduce ConDiSim — a denoising conditional diffusion model [Ho et al., 2020] that systematically addresses the challenge of intractable likelihoods by leveraging samples from the joint distribution $p(\bm{\theta},\mathbf{y})$ , enabling efficient and accurate posterior inference without the need for explicit likelihood evaluations. The model consists of a forward diffusion process that gradually adds Gaussian noise to the parameters $\bm{\theta}$, transforming them into a tractable noise distribution. The reverse process then denoises these parameters, conditioned on observed data $\mathbf{y}$, to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$. Figure 1 provides a schematic overview of the ConDiSim model.
In the forward diffusion process, we construct a discrete-time Markov chain $\{\bm{\theta}_{t}\}_{t=0}^{T}$, initialized at $\bm{\theta}_{0}=\bm{\theta}$. At each step $t$, the state $\bm{\theta}_{t}$ is derived from $\bm{\theta}_{t-1}$ by perturbing it with Gaussian noise according to the transition kernel:
$q(\bm{\theta}_{t}\mid\bm{\theta}_{t-1})=\mathcal{N}\big{(}\bm{\theta}_{t};\,\sqrt{1-\beta_{t}}\,\bm{\theta}_{t-1},\,\beta_{t}\mathbf{I}\big{)},$ (2)
where $\beta_{t}\in(0,1)$ determines the noise level at each step, and $\mathbf{I}$ is the identity matrix. This step-wise transformation progressively corrupts $\bm{\theta}_{0}$, moving its mean towards zero and increasing its variance to unity, such that as $t\to T$, the state converges to $q(\bm{\theta}_{T})=\mathcal{N}(\mathbf{0},\mathbf{I})$. Alternatively, $\bm{\theta}_{t}$ can be directly computed from $\bm{\theta}_{0}$ using the marginal distribution:
$q(\bm{\theta}_{t}\mid\bm{\theta}_{0})=\mathcal{N}\big{(}\bm{\theta}_{t};\,\sqrt{\bar{\alpha}_{t}}\,\bm{\theta}_{0},\,(1-\bar{\alpha}_{t})\mathbf{I}\big{)},$ (3)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s})$ is the cumulative product of the noise scaling factors. This formulation allows efficient computation of $\bm{\theta}_{t}$ at any step $t$, directly linking $\bm{\theta}_{t}$ to $\bm{\theta}_{0}$. The joint forward distribution is given by:
$q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})=q(\bm{\theta}_{T}\mid\bm{\theta}_{0})\prod_{t=2}^{T}q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0}).$ (4)
In the reverse diffusion process, latent states $\{\bm{\theta}_{t}\}_{t=T}^{0}$ are denoised from $\bm{\theta}_{T}\sim\mathcal{N}(\bm{0},\bm{I})$ to recover $\bm{\theta}_{0}$, with every step conditioned on observed data $\mathbf{y}$ [Saharia et al., 2021]. The joint reverse distribution is formulated as:
$p(\bm{\theta}_{0:T}\mid\mathbf{y})=p(\bm{\theta}_{T})\,\prod_{t=1}^{T}p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}),$ (5)
Training the reverse model involves maximizing the log-posterior of the latent variables given the observations:
$\mathcal{L}_{\mathrm{ConDiSim}}=\max\log p(\bm{\theta}_{0}\mid\mathbf{y}),$ (6)
[…]
However, evaluating $\log p(\bm{\theta}_{0}\mid\mathbf{y})$ directly is intractable, as it requires integrating over the full trajectory of intermediate latent states. To address this, we employ a variational approach by introducing the joint forward distribution $q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})$. Using Jensen’s inequality, the lower bound on the log-posterior is given by:
$\log p(\bm{\theta}_{0}\mid\mathbf{y})\;\geq\;\mathbb{E}_{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\!\biggl{[}\log\frac{p(\bm{\theta}_{0:T}\mid\mathbf{y})}{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\biggr{]}.$ (8)
Applying Eqs. (4) and (5) to Eq. (8) leads to the following decomposition:
$\mathcal{L}_{\mathrm{ConDiSim}}=\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{prior}}+\mathcal{L}_{\mathrm{denoise}}$ (9)
where, the reconstruction term, $\mathcal{L}_{\mathrm{recon}}=\mathbb{E}_{q}[\log p(\bm{\theta}_{0}\mid\bm{\theta}_{1},\mathbf{y})]$, evaluates how well the original parameters are recovered from the final denoised state. The prior term, $\mathcal{L}_{\mathrm{prior}}=-D_{\mathrm{KL}}(q(\bm{\theta}_{T}\mid\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{T}))$, enforces consistency with the prior at the final noised step. The denoising term, $\mathcal{L}_{\mathrm{denoise}}=-\sum_{t=2}^{T}\mathbb{E}_{q}[D_{\mathrm{KL}}(q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}))]$, aligns reverse transitions with the forward process. For sufficiently large $T$, the reconstruction and prior terms vanish, leaving the denoising KL as the dominant training objective:
$\mathcal{L}=\sum_{t=2}^{T}\mathbb{E}_{q}\bigl{[}D_{\mathrm{KL}}\bigl{(}q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\;\|\;p_{\phi}(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y})\bigr{)}\bigr{]}.$ (10)
The first term in the KL divergence, $q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})$, can be written using Bayes’ theorem as $q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})=\frac{q(\bm{\theta}_{t}\mid\bm{\theta}_{t-1},\bm{\theta}_{0})\,q(\bm{\theta}_{t-1}\mid\bm{\theta}_{0})}{q(\bm{\theta}_{t}\mid\bm{\theta}_{0})}$. From Eq. (2) and Eq. (3), each term is Gaussian, so the resulting distribution is also Gaussian: $q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})=\mathcal{N}(\bm{\theta}_{t-1};\bm{\mu}_{q},\bm{\Sigma}_{q})$, where:
$\bm{\mu}_{q}$ $=\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\,\bm{\theta}_{0}+\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\,\bm{\theta}_{t},$ (11)
$\bm{\Sigma}_{q}$ $=\frac{(1-\bar{\alpha}_{t-1})\alpha_{t}}{1-\bar{\alpha}_{t}}\,\mathbf{I}.$
To eliminate the explicit dependence on $\bm{\theta}_{0}$, we substitute $\bm{\theta}_{0}=(\bm{\theta}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon})/\sqrt{\bar{\alpha}_{t}}$ (with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$) into Eq. (3), thereby reparameterizing $\bm{\mu}_{q}$ to yield a noise-parameterized expression for the mean:
$\bm{\mu}_{q}(\bm{\theta}_{t},t)$ $=\tfrac{1}{\sqrt{\alpha_{t}}}\Bigl{(}\bm{\theta}_{t}-\tfrac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}\Bigr{)}.$ (12)
Assuming the learned reverse process preserves the covariance structure ($\bm{\Sigma}_{p}=\bm{\Sigma}_{q}$), the KL divergence in Eq. (10) reduces to the squared Euclidean distance between the means:
$\mathcal{L}=\sum_{t=2}^{T}\mathbb{E}_{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\Bigl{[}\big{\lVert}\bm{\mu}_{q}(\bm{\theta}_{t},\bm{\theta}_{0})\;-\;\bm{\mu}_{p}(\bm{\theta}_{t},\mathbf{y},t)\big{\rVert}^{2}\Bigr{]}.$ (13)
From Eq. (12), the mean of learned reverse process can be expressed as:
$\bm{\mu}_{p}(\bm{\theta}_{t},\mathbf{y},t)$ $=\tfrac{1}{\sqrt{\alpha_{t}}}\Bigl{(}\bm{\theta}_{t}-\tfrac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\Bigr{)},$ (14)
where, $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ is neural network that, given the noisy latent $\bm{\theta}_{t}$, conditioning $\mathbf{y}$, and time step $t$, predicts the noise. Substituting the mean parameterizations from Eqs.(12) and (14) into the loss of Eq.(13) and discarding all constant factors, yields the compact noise–prediction objective:
$\mathcal{L}_{\text{ConDiSim}}=\mathbb{E}_{\,t,\,\bm{\epsilon},\,(\bm{\theta}_{0},\mathbf{y})\sim p(\theta,y)}\Bigl{[}\big{\lVert}\bm{\epsilon}-\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\big{\rVert}^{2}\Bigr{]}.$ (15)
Figure 1: The ConDiSim Model: (a) The forward diffusion process incrementally adds Gaussian noise to the parameter vector ($\bm{\theta}$), generating a sequence of latent states ($\{\bm{\theta}_{t}\}$); (b) The reverse diffusion process reconstructs the posterior distribution ($p(\bm{\theta}\mid\mathbf{y})$) by leveraging the observed data ($\mathbf{y}$) for denoising.
3.2 Amortized Inference with ConDiSim
In our framework, we implement the reverse diffusion process as an instance of amortized inference. For a given observation $\mathbf{y_{o}}$, the reverse process is initialized by drawing $N$ latent variables $\bm{\theta}_{T}^{(i)}$ from the standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ for $i=1,\dots,N$. These latent samples are then conditioned on $\mathbf{y_{o}}$ and iteratively refined through the learned reverse diffusion process. Specifically, at each reverse time step, the transition is characterized by,
$\bm{\theta}_{t-1}^{(i)}\mid\bm{\theta}_{t}^{(i)},\bm{y_{o}}\sim\mathcal{N}\!\left(\bm{\mu}_{\phi}\!\left(\bm{\theta}_{t}^{(i)},\mathbf{y_{o}},t\right),\,\bm{\Sigma}_{t}\right),$
where $\bm{\mu}_{\phi}$ and $\bm{\Sigma}_{t}$ represent the learned mean and covariance parameters that determine the probabilistic mapping between consecutive latent states in the reverse process. After iterating from $t=T$ down to $t=0$, the process yields refined latent samples $\bm{\theta}_{0}^{(i)}$ that approximate the target posterior distribution $p(\bm{\theta}\mid\mathbf{y_{o}})$. This method efficiently produces high-quality posterior samples suitable for applications such as parameter estimation and uncertainty quantification.
4 Experiments
We evaluate ConDiSim on 10 benchmark problems from (sbibm) suite [Lueckmann et al., 2021] and two real-world problems. For each benchmark, we train with simulation budgets of 10 0001000010\,00010 000, 20 0002000020\,00020 000, and 30 0003000030\,00030 000 simulations, and infer the posterior by drawing 10 0001000010\,00010 000 samples from the model. Posterior quality is assessed with Maximum Mean Discrepancy (MMD), the Classifier Two-Sample Test (C2ST), and Empirical Cumulative Distribution Function (ECDF) diagnostics implemented in BayesFlow [Radev et al., 2023b]. ECDF comparisons were performed using the BayesFlow framework [Radev et al., 2023b]; further details and ECDF plots for various benchmarks are provided in Appendix E. Complete hyperparameter configurations for all methods are detailed in Appendix B.
(a) Posterior Plot
(b) ECDF Plot
Figure 2: Two Moons: Comparison of generated and reference posterior distributions, with the ECDF plot representing the cumulative distribution of fractional rank statistics from posterior draws.
(a) SLCP
(b) SLCP Distractors
Figure 3: SLCP: Comparison of generated vs. reference posteriors for SLCP and SLCP Distractors
4.1 Benchmark Posterior Analysis
In this section, we analyze three representative benchmarks—Two Moons, SLCP, and SLCP Distractors—selected for their distinct challenges in posterior inference. The Two Moons benchmark poses the challenge of capturing a bi-modal posterior with a distinctive crescent geometry. As depicted in Figure 2(a), ConDiSim accurately recovers both modes and the complex posterior structure. The ECDF plots for $\theta_{1}$ and $\theta_{2}$ (Figure 2(b)) demonstrate that the inferred distributions are well-calibrated, with the majority of rank statistics falling within the 90% confidence bands; a minor deviation for $\theta_{1}$ suggests a slight degree of overconfidence.
The SLCP benchmark encompasses a range of parameter behaviors - including uniform, Gaussian, multi-modal, and skewed distributions. Its variant, SLCP Distractors, further tests the model’s robustness by incorporating uninformative noise dimensions, thereby simulating scenarios with heavily corrupted data. Figures 3(a) and 3(b) illustrate that ConDiSim is able to capture these complex posterior landscapes with high fidelity, and also effectively isolate the informative parameters from the distractors. The rank ECDF plots for both SLCP variants (Figures 14(a) and 14(b) in the Appendix) also remain centered within the confidence intervals, confirming that the model reliably captures parameter variability even under noisy conditions. Detailed descriptions of these benchmarks, are provided in Appendix A.
Table 1 presents the C2ST scores for the 10 benchmarking problems. For the Two Moons task, ConDiSim achieves C2ST performance comparable to Simformer, APT, and NPE, while requiring the shortest training time among the evaluated models (see Table 2). Simformer and APT perform well on most test problems – Simformer leveraging its transformer architecture with cross-attention, while APT is able to iteratively refine the posterior approximation. The SLCP problem remains particularly challenging, with all methods struggling to accurately infer the true posterior. It is important to note that the C2ST metric is sensitive to parameter estimation errors and can be difficult to interpret in high-dimensional settings, as a single poorly estimated dimension may disproportionately affect the overall score. For instance, as illustrated in Figure 3(a), ConDiSim encounters difficulties with the final dimension, despite accurately capturing the other dimensions, which is reflected in an overall C2ST score of 0.82. To provide additional insight, we report the MMD metric for each benchmark problem in the Supplementary Table 4.
(a) Hodgkin–Huxley: Approximated Posterior
(b) Posterior Predictive Voltage Traces
Figure 4: Posterior Distributions for Hodgkin–Huxley model
(a) Genetic Oscillator: Approximated Posterior
(b) Posterior Predictive Trajectories
Figure 5: Posterior Distributions for Genetic Oscillator model
4.2 Application to Real-World Scenarios
We evaluate the performance of ConDiSim on two real-world simulation models. First, the Hodgkin-Huxley (HH) model [Gloeckler et al., 2024] which simulates neuronal action potentials by modeling the membrane voltage dynamics governed by sodium, potassium, and leak currents. Second, we consider the Genetic Oscillator model which is a high-dimensional gene regulatory network that exhibits oscillatory behavior, driven by a positive-negative feedback loop mimicking a circadian clock [Vilar et al., 2002]. The model encompasses 9 species undergoing 18 reactions, parameterized by 15 reaction rate constants. Although the full system involves multiple interacting components, we focus on the inference of only three species (C, A, R). Detailed description is available in Appendix A.
Unlike the SBI benchmarking problems, the true posterior distributions for these real-world models are unknown. To assess ConDiSim, we use well motivated parameter combinations from literature to generate corresponding synthetic observations (e.g., parameter combination resulting in stable oscillations in case of the genetic oscillator). These observations serve as the observed data for inference, enabling us to evaluate the inferred posterior by comparing its concentration around the true parameter value. We also perform a posterior predictive check to validate the model. In this procedure, five posterior samples are randomly selected and used to simulate corresponding observations, which are then compared with the observed data. This check provides an additional measure of the consistency and reliability of the inferred posteriors.
For the HH model, Figure 4(a) shows that ConDiSim effectively infers a posterior distribution concentrated around the true parameter values, while Figure 4(b) confirms that the simulated voltage trajectories closely match the reference signals. The parameter estimation task for the genetic oscillator is more challenging due to broad and varied parameter priors. Figure 5(a) indicates that the peak of the posterior distribution is near the true parameter values. Figure 5(b) illustrates both the true and generated trajectories for three selected species, obtained by randomly sampling three posterior sets and plotting the corresponding simulations. Initially, these trajectories align well with the true data, with slight deviations emerging over time due to the inherent stochastic nature of the model. Figure 7 (Appendix) depicts multiple trajectories simulated using the true parameter combination - which closely match the inferred results. In both cases, the trajectories are very close to each other during initial time steps, and develop small deviations owing to model stochasticity as time steps evolve. The plots highlight the system’s inherent variability, and also validating the inferred solution.
5 Discussion
We analyze the impact of the hyperparameters — variance schedule ($\beta_{t}$) and diffusion steps ($T$), on ConDiSim’s performance in Section 5.1, with further discussion on overall scalability, robustness, and reliability in Sections 5.2, 5.4, and 5.5 respectively. This study provides practical insights for optimizing accuracy, efficiency, and reliability across various simulation-based inference tasks.
5.1 Variance Schedule & Diffusion Steps
We conducted an ablation study on the Two Moons test problem to examine the effects of diffusion steps ($T$) and variance schedules ($\beta_{t}$) on posterior approximation accuracy. Varying $T$ from 100 to 1000, we calculate the mean Classifier Two-Sample Test (C2ST) accuracy across five runs with a fixed simulation budget of 10,000 samples, while holding architecture and other hyperparameters constant. The study compared linear, cosine, and quadratic schedules to assess their impact under varying diffusion granularity, with $T$ controlling the refinement steps and $\beta_{t}$ setting the noise rate in the forward diffusion process.
Figure 6: Mean C2ST score across diffusion steps for linear, cosine, and quadratic schedules on the Two Moons benchmark.
Table 1: Comparison of C2ST scores across SBI benchmarks: (mean $\pm$ std.) at varying simulation budgets.
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 1.0 $\pm$0.0000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0000 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.$9$$\pm$0.0003 0.$9$$\pm$0.0012 1.$0$$\pm$0.0002 FMPE 0.9713 $\pm$ 0.0117 0.9779 $\pm$ 0.0095 0.9387 $\pm$ 0.0377 0.9742 $\pm$ 0.0113 0.6661 $\pm$ 0.0307 0.6653 $\pm$ 0.0377 0.$9$$\pm$ 0.0178 0.$9$$\pm$ 0.0345 0.$9$$\pm$ 0.0018 ConDiSim 0.9722 $\pm$0.0368 0.9366 $\pm$0.02395 0.9347 $\pm$0.0368 0.6430 $\pm$0.0237 0.6155 $\pm$0.0317 0.6166 $\pm$0.0480 0.9996 $\pm$0.0002 0.9995 $\pm$0.0002 0.9996 $\pm$0.0002
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7474 $\pm$0.0364 0.6977 $\pm$0.0260 0.7234 $\pm$0.0599 0.9930 $\pm$0.0038 0.9916 $\pm$0.0046 0.9929 $\pm$0.0059 0.9989 $\pm$0.0004 0.9976 $\pm$0.0016 0.9988 $\pm$0.0010 0.9687 $\pm$0.0286 0.9436 $\pm$0.0203 0.9723 $\pm$0.0140 NPE 0.6118 $\pm$0.0155 0.5830 $\pm$0.0240 0.5818 $\pm$0.0158 0.5472 $\pm$0.0076 0.5238 $\pm$0.0112 0.5266 $\pm$0.0037 0.5562 $\pm$0.0147 0.5292 $\pm$0.0068 0.5256 $\pm$0.0066 0.8996 $\pm$0.0206 0.8502 $\pm$0.0174 0.8546 $\pm$0.0213 APT 0.5457 $\pm$0.0058 0.5425 $\pm$0.0124 0.5467 $\pm$0.0050 0.5367 $\pm$0.0038 0.5293 $\pm$0.0052 0.5258 $\pm$0.0047 0.5401 $\pm$0.0115 0.5302 $\pm$0.0069 0.5264 $\pm$0.0075 0.8164 $\pm$0.0295 0.6882 $\pm$0.0342 0.6473 $\pm$0.0343 Simformer 0.5166 $\pm$0.0210 0.5100 $\pm$0.0090 0.5126 $\pm$0.0070 0.6130 $\pm$0.0281 0.6226 $\pm$0.0398 0.5866 $\pm$0.0430 0.6501 $\pm$0.0519 0.6207 $\pm$0.0519 0.5780 $\pm$0.0211 0.7661 $\pm$0.0685 0.6201 $\pm$0.0490 0.7192 $\pm$0.0721 FMPE 0.7894 $\pm$ 0.0254 0.7872 $\pm$ 0.0049 0.5821 $\pm$ 0.0292 0.9605$\pm$ 0.0019 0.9716 $\pm$ 0.0023 0.9708 $\pm$ 0.0025 0.9218 $\pm$ 0.0122 0.9100 $\pm$ 0.0174 0.9082 $\pm$ 0.0178 0.9698 $\pm$ 0.0132 0.9686 $\pm$ 0.0139 0.9686 $\pm$ 0.0139 ConDiSim 0.6602 $\pm$0.0098 0.6181 $\pm$0.0237 0.6233 $\pm$0.0276 0.5809 $\pm$0.0210 0.5656 $\pm$0.0197 0.5505 $\pm$0.0056 0.6572 $\pm$0.0215 0.6468 $\pm$0.0364 0.6448 $\pm$0.0216 0.8691 $\pm$0.0132 0.8324 $\pm$0.0281 0.8210 $\pm$0.0188
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 $\pm$0.0630 0.6634 $\pm$0.0503 0.7467 $\pm$0.1036 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.9997 $\pm$0.0005 0.9990 $\pm$0.0015 0.9912 $\pm$0.0161 0.9838 $\pm$0.0157 NPE 0.5749 $\pm$0.0331 0.5931 $\pm$0.0329 0.5411 $\pm$0.0195 0.9121 $\pm$0.0338 0.8988 $\pm$0.0306 0.8898 $\pm$0.0565 0.8378 $\pm$0.0396 0.9122 $\pm$0.0207 0.9405 $\pm$0.0125 APT 0.5189 $\pm$0.0080 0.5202 $\pm$0.0114 0.5147 $\pm$0.0125 0.8404 $\pm$0.0107 0.8328 $\pm$0.0156 0.8353 $\pm$0.0164 0.9433 $\pm$0.0076 0.9434 $\pm$0.0061 0.9457 $\pm$0.0068 Simformer 0.5273 $\pm$0.0229 0.5190 $\pm$0.0174 0.5098 $\pm$0.0084 0.6584 $\pm$0.0381 0.6131 $\pm$0.0225 0.6154 $\pm$0.0265 0.9378 $\pm$0.0053 0.9520 $\pm$0.0064 0.9478 $\pm$0.0049 FMPE 0.8051 $\pm$ 0.0506 0.7745 $\pm$ 0.0261 0.5706 $\pm$ 0.0209 0.9266 $\pm$ 0.0288 0.9118 $\pm$ 0.0376 0.8431 $\pm$ 0.0393 0.9573 $\pm$ 0.0013 0.8454 $\pm$ 0.0014 0.8101 $\pm$ 0.0020 ConDiSim 0.5291 $\pm$0.0143 0.5395 $\pm$0.0248 0.5242 $\pm$0.0063 0.6589 $\pm$0.0421 0.5892 $\pm$0.0100 0.6146 $\pm$0.0371 0.8324 $\pm$0.0332 0.8083 $\pm$0.0157 0.8043 $\pm$0.0138
Table 2: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.).
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 $\pm$2.43 164.90 $\pm$57.10 249.50 $\pm$105.05 118.62 $\pm$0.18 179.69 $\pm$49.96 253.48 $\pm$113.38 227.16 $\pm$1.44 241.24 $\pm$7.58 250.47 $\pm$2.22 NPE 7.04 $\pm$0.21 14.21 $\pm$0.04 20.95 $\pm$0.37 12.17 $\pm$0.17 24.96 $\pm$0.64 37.43 $\pm$0.22 20.10 $\pm$5.45 29.62 $\pm$3.74 44.92 $\pm$9.24 APT 90.31 $\pm$8.99 186.17 $\pm$13.79 268.28 $\pm$32.88 101.39 $\pm$7.89 197.30 $\pm$9.32 267.38 $\pm$21.93 40.15 $\pm$3.90 73.74 $\pm$6.25 112.51 $\pm$2.51 Simformer 3.96 $\pm$0.04 7.36 $\pm$0.11 10.73 $\pm$0.25 17.87 $\pm$0.18 35.05 $\pm$0.52 52.81 $\pm$0.91 4.24 $\pm$0.29 7.67 $\pm$0.38 10.68 $\pm$0.45 FMPE 2.63 $\pm$ 0.05 4.73$\pm$ 1.11 10.655 $\pm$ 0.04 2.65 $\pm$ 0.05 5.75 $\pm$ 1.43 7.64 $\pm$ 1.38 2.78 $\pm$ 0.15 6.39 $\pm$ 0.19 8.94 $\pm$ 2.94 CondiSim 0.84 $\pm$0.28 1.682 $\pm$0.09 2.835 $\pm$0.70 0.31 $\pm$0.11 0.60 $\pm$0.14 0.88 $\pm$0.17 0.87 $\pm$0.16 0.94 $\pm$0.13 1.59 $\pm$0.24
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 FMPE 2.76 $\pm$ 0.08 3.08 $\pm$ 1.76 12.97 $\pm$ 0.87 2.44 $\pm$ 0.44 4.42 $\pm$ 1.37 8.05 $\pm$ 1.38 3.63 $\pm$ 0.42 6.97 $\pm$ 0.03 9.55 $\pm$ 0.72 CondiSim 1.31 $\pm$0.07 1.83 $\pm$0.32 2.08 $\pm$0.34 0.43 $\pm$0.04 0.62 $\pm$0.08 0.93 $\pm$0.17 0.60 $\pm$0.11 1.15 $\pm$0.04 1.41 $\pm$0.08
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 114.07 $\pm$2.11 111.30 $\pm$0.28 137.46 $\pm$19.16 48.19 $\pm$0.24 48.70 $\pm$0.84 49.12 $\pm$0.49 43.27 $\pm$0.36 66.43 $\pm$5.70 69.10 $\pm$0.15 162.72 $\pm$14.73 196.19 $\pm$119.42 99.39 $\pm$5.63 NPE 19.56 $\pm$2.47 42.88 $\pm$4.54 64.03 $\pm$6.51 12.03 $\pm$0.80 24.69 $\pm$1.63 38.79 $\pm$1.23 16.55 $\pm$0.85 30.90 $\pm$2.77 49.42 $\pm$1.49 42.97 $\pm$6.35 97.70 $\pm$6.66 117.61 $\pm$4.12 APT 76.27 $\pm$7.02 167.88 $\pm$15.99 240.13 $\pm$22.74 70.47 $\pm$0.88 135.38 $\pm$5.75 203.27 $\pm$5.97 78.61 $\pm$2.11 158.57 $\pm$9.32 219.44 $\pm$11.83 93.40 $\pm$1.26 150.29 $\pm$11.10 216.30 $\pm$21.38 Simformer 3.84 $\pm$0.04 7.15 $\pm$0.09 10.48 $\pm$0.21 18.18 $\pm$0.13 35.46 $\pm$0.17 55.46 $\pm$1.13 17.89 $\pm$0.23 33.47 $\pm$0.40 51.24 $\pm$0.89 4.23 $\pm$0.01 7.67 $\pm$0.13 11.32 $\pm$0.40 FMPE 2.16$\pm$ 0.08 5.36 $\pm$ 0.78 8.31 $\pm$ 1.32 1.78 $\pm$ 0.05 4.40 $\pm$ 0.78 5.94 $\pm$ 4.31 2.15 $\pm$ 0.22 4.30 $\pm$ 3.86 8.71 $\pm$ 2.12 1.36 $\pm$ 0.14 3.71 $\pm$ 2.57 6.34 $\pm$ 2.33 ConDiSim 0.74 $\pm$0.18 1.77 $\pm$0.41 2.31 $\pm$0.59 0.76 $\pm$0.09 1.95 $\pm$0.59 3.46 $\pm$0.70 0.91 $\pm$0.20 2.34 $\pm$0.37 3.02 $\pm$0.69 1.13 $\pm$0.08 2.61 $\pm$0.80 2.82 $\pm$0.36
As shown in Figure 6, lower $T$ values generally yield better C2ST accuracy, potentially due to reduced cumulative noise, which preserves crucial posterior structure. The quadratic schedule outperforms others, possibly because its gradual noise addition captures complex structures more effectively without early-stage noise overload. Increasing $T$ leads to a plateau and slight decline in accuracy, suggesting diminishing returns and rising instability, with all schedules converging in performance at high $T$ values.
5.2 Scalability
ConDiSim’s scalability was extensively evaluated across diverse benchmark problems, spanning from low-dimensional (Two Moons, GMM, SIR; 2 parameters) to higher-dimensional scenarios (Gaussian Linear, Bernoulli GLM; 10 parameters, HH Model; 7 parameters, genetic oscillator; 15 parameters). From the C2ST metric values in Table 1, and the posterior plots spanning from Figure 2(a) to Figure 13, it can be seen that ConDiSim maintained the posterior approximation accuracy with increasing problem dimensionality and complexity. Model training times in Table 2 also show that ConDiSim is computationally efficient to train and scales favorably with problem complexity, making it suitable for high-dimensional SBI tasks. Its ability to handle larger parameter spaces without significant loss of performance is a key advantage, as evidenced in the genetic oscillator test problem involving 15 parameters.
5.3 Comparative Evaluation of different Inference Methods
Tables 1 and 4 indicate that diffusion models (Simformer, ConDiSim) generally outperform flow-based (NPE, APT) and GAN-based (GATSBI) approaches in accuracy, owing to their ability to model complex posteriors. Simformer achieves more accurate posterior approximations in most test problems (also see Table 4) owing to its complex architecture, however it also entails longer training times due to the involved complexity (Table 2). In contrast, ConDiSim’s simplified conditioning mechanism and architecture reduces training complexity, and its iterative denoising process does not impede inference speed in our experiments. The known challenges of adversarial training affect GATSBI, while APT’s sequential nature increases its computational cost relative to NPE.
5.4 Robustness
Robustness to observational noise is a critical factor in the practical application of simulation-based inference (SBI) methods. To rigorously evaluate this aspect, we conducted experiments on two challenging tasks: the SLCP Distractors and the Bernoulli GLM Raw. The SLCP Distractors task introduces 92 non-informative variables to the observations, significantly increasing the inference complexity. In contrast, the Bernoulli GLM Raw task utilizes raw data instead of well-behaved summary statistics, presenting a more realistic scenario often encountered in real-world applications. The results for the SLCP Distractors and Bernoulli GLM RAW, illustrated in Table 1 and Figures 3(b),12(b) demonstrate ConDiSim’s robustness to both observational noise. Such resilience is important as it indicates ConDiSim’s capacity to effectively handle real-world data, which is often characterized by measurement errors and inherent uncertainties.
5.5 Reliability
To further validate the learned inference model, we employed Simulation-Based Calibration (SBC) as a rigorous diagnostic for assessing the quality of approximate posterior distributions. The empirical cumulative distribution function (ECDF) plots demonstrate ConDiSim’s ability to produce accurate and well-calibrated posteriors, with rank statistics closely aligning along the diagonal—indicative of unbiased and reliable inference. This alignment supports the statistical validity and robustness of ConDiSim’s inferences across diverse and challenging scenarios.
6 Conclusion
In this work, we introduced ConDiSim, a denoising conditional diffusion model for simulation-based inference. We demonstrated ConDiSim’s ability to effectively capture complex posterior distributions efficiently on 10 benchmark problems and 2 real-world test problems. ConDiSim performs comparably to the state-of-the-art while being substantially faster to train. As part of an ablation study, we also explored the significance of the variance schedule and the number of diffusion steps in influencing model performance. ConDiSim offers a flexible and efficient approach to simulation-based inference, with potential applications across various scientific and engineering domains. Future work includes exploring adaptive variance schedules, integrating domain-specific knowledge into the model, and extending ConDiSim towards learning both the likelihood and posterior distributions.
Acknowledgments
The computations/data handling were enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre and by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons at Chalmers, and Uppsala Multidisciplinary Center for Advanced Computational Science (UPPMAX) at Uppsala University, partially funded by the Swedish Research Council through grant agreement nos. 2022-06725 and 2018-05973. Andreas Hellander and Prashant Singh acknowledge support from the Swedish Research Council through grant agreement nos. 2023-05167 and 2023-05593 respectively.
Appendix A SBI Problem Description
We present the general setup for simulation-based inference (SBI), which is employed across various benchmarking problems to estimate the posterior distribution $p(\bm{\theta}|\mathbf{y})$ over parameters $\bm{\theta}$, given observed data $\mathbf{y}$. As outlined in [Lueckmann et al., 2021], the framework for SBI problems typically includes the following components:
  • Prior Distribution: A prior distribution $p(\bm{\theta})$ that defines the initial beliefs about the parameter vector before any data is observed. This distribution serves as the starting point for the inference process.
  • Simulation Model: A stochastic simulation model $\mathcal{M}$ that allows empirical evaluation of the likelihood function $p(\mathbf{y}|\bm{\theta})$, illustrating how synthetic data $\mathbf{y}$ is generated given specific parameter values $\bm{\theta}$. This model encapsulates the underlying mechanisms of the problem and provides samples from the joint distribution $p(\bm{\theta},\mathbf{y})$, enabling inference without direct likelihood computation.
  • Dimensionality: The dimensionality of both the parameter vector $\bm{\theta}$ and the observed data $\mathbf{y}$ is crucial for parameter inference. It can vary greatly between problems, and directly impacts the complexity and computational cost of the inference task. High-dimensional spaces often present greater challenges in accurately estimating the posterior distribution.
A.1 Two Moons
This parameter inference task is designed to test the performance of inference methods under multi-modal conditions, featuring a posterior distribution with both global (bi-modal) and local (crescent-shaped) characteristics.
  • Prior: $\bm{\theta}\sim\mathcal{U}(-1,1)$
  • Simulator: $\mathbf{y}|\bm{\theta}=r\begin{bmatrix}\cos(\alpha)\\ \sin(\alpha)\end{bmatrix}+0.25\begin{bmatrix}-|\theta_{1}+\theta_{2}|/\sqrt{2}\\ (-\theta_{1}+\theta_{2})/\sqrt{2}\end{bmatrix},$ where $\alpha\sim\mathcal{U}(-\pi/2,\pi/2)$ and $r\sim\mathcal{N}(0.1,0.01^{2})$.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
A.2 Gaussian Mixture
This task challenges inference methods to estimate the shared mean of a mixture of Gaussian distributions with distinct covariances, thereby assessing their performance in handling multi-modal and overlapping data distributions.
  • Prior: $\bm{\theta}\sim\mathcal{U}(-10,10)$
  • Simulator: $\mathbf{y}|\bm{\theta}=0.5\left(\mathcal{N}(\bm{\theta},\mathbf{I})+\mathcal{N}(\bm{\theta},0.01\mathbf{I})\right)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
A.3 Gaussian Linear
This task involves estimating the mean vector $\bm{\theta}$ of a multivariate Gaussian distribution with a fixed and known covariance matrix. Since both the prior and the likelihood are Gaussian, they form a conjugate pair, resulting in a Gaussian posterior distribution. This conjugacy makes the task an ideal benchmark for evaluating inference methods, as it allows for a clear assessment of their ability to accurately capture known posterior distributions.
  • Prior: $\bm{\theta}\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$
  • Simulator: Generates data $\mathbf{y}$ using $\mathbf{y}|\bm{\theta}\sim\mathcal{N}(\bm{\theta},0.1\mathbf{I})$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
A.4 Gaussian Linear Uniform
This task is a variant of the Gaussian Linear problem, where the mean parameter $\bm{\theta}$ is assigned a uniform prior distribution $\bm{\theta}\sim\mathcal{U}(-1,1)$ instead of a Gaussian prior. The introduction of a uniform prior imposes bounded support on the parameters, altering the posterior distribution from a Gaussian to a truncated Gaussian. This modification tests the inference method’s ability to handle non-Gaussian priors and to accurately capture the resulting posterior distributions.
A.5 SLCP
The SLCP task presents a more challenging inference scenario where, despite the simplicity of the likelihood function, the posterior distribution exhibits complex, non-linear dependencies among parameters. The parameter vector $\bm{\theta}$ is five-dimensional, and the simulator generates two-dimensional data points by applying non-linear transformations to $\bm{\theta}$. This setup assesses the inference method’s capability to accurately capture intricate posterior structures arising from non-linear relationships between parameters and observations.
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
  • Simulator: Outputs data $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{4})$, where:
    $y_{i}\sim\mathcal{N}\left(\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},\begin{bmatrix}\theta_{3}^{2}&\tanh(\theta_{5})\theta_{3}\theta_{4}\\ \tanh(\theta_{5})\theta_{3}\theta_{4}&\theta_{4}^{2}\end{bmatrix}\right)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{8}$
A.6 SLCP with Distractors
An extension of the SLCP task that incorporates additional non-informative dimensions (distractors) into the observed data. This setup evaluates the robustness of inference methods when faced with high-dimensional data containing redundant information.
  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
  • Simulator: Outputs data $y|\theta=(y_{1},\ldots,y_{100})$, where a permutation function $p(y)$ reorders $y$ with a fixed random permutation.
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{100}$
A.7 Bernoulli GLM
This task involves performing inference in a Generalized Linear Model (GLM) with Bernoulli-distributed observations and a 10-dimensional parameter vector $\bm{\theta}$. In this model, each binary observation is linked to the parameters through a logistic function, modeling the probability of success in a Bernoulli trial. Gaussian priors are imposed on the coefficients $\beta$ and the function f, promoting smoothness in the parameter estimates. This setup evaluates the inference method’s ability to handle discrete data and accurately capture the posterior distribution in models with latent smoothness constraints.
  • Prior: $\beta\sim\mathcal{N}(\mathbf{0},2)$, $f\sim\mathcal{N}(\mathbf{0},(\mathbf{F}^{\top}\mathbf{F})^{-1})$
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$, with $\eta(z)=\frac{\exp(z)}{1+\exp(z)}$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
  • Fixed Parameters: $T=100$
A.8 Bernoulli GLM Raw
This task is similar to the Bernoulli GLM described above but utilizes raw binary observations instead of sufficient statistics. Without the simplifying benefit of sufficient statistics, the inference method must directly model the relationship between the parameters and the raw data, increasing the complexity of the inference task. This scenario tests the method’s capability to accurately infer parameters in the presence of high-dimensional, discrete, and unprocessed data, challenging its robustness and flexibility.
  • Prior: Same as Bernoulli GLM
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{100}$
  • Fixed Parameters: $T=100$
A.9 SIR Model
The SIR model is an epidemiological framework that describes the number of individuals in three compartments: Susceptible ($S$), Infectious ($I$), and Recovered ($R$). The dynamics of the system are governed by the following differential equations:
$\frac{d}{dt}\begin{pmatrix}S\\ I\\ R\end{pmatrix}=\begin{pmatrix}-\beta S\\ \beta S-\gamma I\\ \gamma I\end{pmatrix}\frac{I}{N},$ (16)
where $N$ represents the total population size. The parameter vector $\bm{\theta}$ is defined as:
$\bm{\theta}=\begin{pmatrix}\beta\\ \gamma\end{pmatrix}\sim\text{LogNormal}\left(\begin{pmatrix}\log(0.4)\\ \log(1/8)\end{pmatrix},\begin{pmatrix}0.5\\ 0.2\end{pmatrix}\right).$
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\mathcal{B}(1000,\frac{I}{N})$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{10}$
  • Fixed Parameters: Population size $N=1,000,000$, task duration $T=160$
A.10 Lotka-Volterra
The Lotka-Volterra (LV) model is a classical ecological framework that describes the interaction between prey ($X$) and predator ($Y$) populations. The dynamics are expressed as:
$\frac{d}{dt}\begin{pmatrix}X\\ Y\end{pmatrix}=\begin{pmatrix}\alpha X-\beta XY\\ -\gamma Y+\delta XY\end{pmatrix},$ (17)
with the parameter vector defined as:
$\bm{\theta}=\begin{pmatrix}\alpha\\ \beta\\ \gamma\\ \delta\end{pmatrix}\sim\text{LogNormal}\left(\begin{pmatrix}-0.125\\ -3\\ -0.125\\ -3\end{pmatrix},\begin{pmatrix}0.5\\ 0.5\\ 0.5\\ 0.5\end{pmatrix}\right).$
  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\text{LogNormal}(\log(X),0.1)$ and $y_{2i}\sim\text{LogNormal}(\log(Y),0.1)$
  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{4},\mathbf{y}\in\mathbb{R}^{20}$
  • Fixed Parameters: Task duration $T=20$, initial conditions $(X(0),Y(0))=(30,1)$
A.11 Hodgkin-Huxley Model
The Hodgkin-Huxley (HH) model describes the generation and propagation of action potentials in neurons by modeling the membrane voltage $V(t)$ as influenced by ionic currents through voltage-gated sodium ($\text{Na}^{+}$), potassium ($\text{K}^{+}$), and leak channels. We adopt the formulation from [Gloeckler et al., 2024]. The model is given by the following set of stochastic differential equations:
$\frac{dV}{dt}=\frac{I_{\text{inj}}(t)-g_{\text{Na}}m^{3}h(V-E_{\text{Na}})-g_{\text{K}}n^{4}(V-E_{\text{K}})-g_{\text{L}}(V-E_{\text{L}})}{C_{m}}+0.05\,dw_{t},$ (18)
where $dw_{t}$ represents a Gaussian white noise term that captures stochastic fluctuations in the membrane potential.
The gating variables $m(t),h(t),n(t)$ follow first-order kinetics:
$\frac{dm}{dt}$ $=\alpha_{m}(V)(1-m)-\beta_{m}(V)m,$ (19)
$\frac{dh}{dt}$ $=\alpha_{h}(V)(1-h)-\beta_{h}(V)h,$ (20)
$\frac{dn}{dt}$ $=\alpha_{n}(V)(1-n)-\beta_{n}(V)n,$ (21)
The voltage-dependent rate functions $\alpha_{x}(V)$ and $\beta_{x}(V)(x\in\{m,h,n\})$ defined as follows:
$\alpha_{m}(V)=0.32\times\frac{\text{efun}(-0.25(V-V_{0}-13.0))}{0.25},\quad\beta_{m}(V)=0.28\times\frac{\text{efun}(0.2(V-V_{0}-40.0))}{0.2},$
$\alpha_{h}(V)=0.128\times\exp\left(\frac{-(V-V_{0}-17.0)}{18.0}\right),\quad\beta_{h}(V)=\frac{4.0}{1+\exp\left(\frac{-(V-V_{0}-40.0)}{5.0}\right)},$
$\alpha_{n}(V)=0.032\times\frac{\text{efun}(-0.2(V-V_{0}-15.0))}{0.2},\quad\beta_{n}(V)=0.5\times\exp\left(\frac{-(V-V_{0}-10.0)}{40.0}\right).$
Here, $\text{efun}(x)$ is defined to ensure numerical stability when evaluating small arguments:
$\text{efun}(x)=\begin{cases}1-\frac{x}{2}&\text{if }x<1\text{e-4},\\ \frac{x}{\exp(x)-1.0}&\text{otherwise}.\end{cases}$
The metabolic cost $H(t)$, representing the energy consumption due to sodium channel activity, is modeled as:
$\frac{dH}{dt}=g_{\text{Na}}m^{3}h(V-E_{\text{Na}}).$ (22)
  • Parameters: The parameter vector is $\bm{\theta}=(C_{m},g_{\text{Na}},g_{\text{K}},g_{\text{L}},E_{\text{Na}},E_{\text{K}},E_{\text{L}})$, where $C_{m}$ is the membrane capacitance, $g_{\text{Na}},g_{\text{K}},g_{\text{L}}$ are the conductances for sodium, potassium, and leak currents, and $E_{\text{Na}},E_{\text{K}},E_{\text{L}}$ are the corresponding reversal potentials.
  • Prior: The parameters $\bm{\theta}$ are sampled from a uniform distribution:
    $\bm{\theta}\sim\mathcal{U}\left([1.0,60,10,0.1,40,-100,-90],[2.0,120,30,0.5,70,-60,-60]\right).$
  • Simulator: The HH model is simulated for 200 ms. An input current of $I_{\text{inj}}(t)=4\,\text{mA}$ is applied between $50$ ms and $150$ ms. The voltage equation is integrated with added Gaussian white noise $(0.05\,dw_{t})$. To reduce dimensionality, the raw voltage trace is summarized into a set of statistics:
    • Spike count,
    • Mean and standard deviation of the resting potential,
    • Mean voltage during spiking,
    • Higher-order moments (normalized skewness and kurtosis) of the voltage response,
    • Total energy consumption due to sodium current.
  • Dimensionality: The parameter space is $\bm{\theta}\in\mathbb{R}^{7}$, and the output summary statistics $\mathbf{y}\in\mathbb{R}^{8}$.
[…]
A.12 High-Dimensional Genetic Oscillator
The genetic oscillator model describes a biochemical reaction network involving $9$ species governed by $18$ reactions, parameterized by $15$ reaction rate constants. This network is based on a positive–negative feedback loop, where the activator protein $A$ upregulates transcription of both its gene and the repressor protein $R$. The repressor $R$ sequesters $A$ by forming a complex $C$, thereby inhibiting further activation [Vilar et al., 2002]. The species in the network are:
  • $D_{A},D_{A}^{*}$: Activator gene in unbound ($D_{A}$) and bound states ($D_{A}^{*}$).
  • $D_{R},D_{R}^{*}$: Repressor gene in unbound ($D_{R}$) and bound states ($D_{R}^{*}$).
  • $M_{A},M_{R}$: mRNA transcribed from activator ($M_{A}$) and repressor genes ($M_{R}$).
  • $A$: Activator protein.
  • $R$: Repressor protein.
  • $C$: Complex formed by $A$ and $R$.
The network includes the following key reactions:
$D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},$ $D_{R}^{*}$ $\xrightarrow{\theta_{R}}D_{R},$
$D_{A},A$ $\xrightarrow{\gamma_{A}}D_{A}^{*},$ $D_{R},A$ $\xrightarrow{\gamma_{R}}D_{R}^{*},$
$D_{A}^{*}$ $\xrightarrow{\alpha_{a}^{\prime}}M_{A},$ $D_{R}^{*}$ $\xrightarrow{\alpha_{r}^{\prime}}M_{R},$
$M_{A}$ $\xrightarrow{\delta_{ma}}\phi,$ $M_{R}$ $\xrightarrow{\delta_{mr}}\phi,$
$M_{A}$ $\xrightarrow{\beta_{a}}M_{A}+A,$ $M_{R}$ $\xrightarrow{\beta_{r}}M_{R}+R,$
$A,R$ $\xrightarrow{\gamma_{c}}C,$ $C$ $\xrightarrow{\delta_{a}}R,$
$A$ $\xrightarrow{\delta_{a}}\phi,$ $R$ $\xrightarrow{\delta_{r}}\phi.$
The reaction rates include:
$\theta_{A},\theta_{R}$ $:\text{Gene unbinding rates},\quad\gamma_{A},\gamma_{R},\gamma_{C}:\text{Binding rates},$
$\alpha_{a},\alpha_{a}^{\prime},\alpha_{r},\alpha_{r}^{\prime}$ $:\text{Transcription rates},\quad\beta_{a},\beta_{r}:\text{Translation rates},$
$\delta_{ma},\delta_{mr}$ $:\text{mRNA degradation rates},\quad\delta_{a},\delta_{r}:\text{Protein degradation rates}.$
Prior:
The parameter vector $\bm{\theta}\in\mathbb{R}^{15}$ is defined with prior bounds:
$\bm{\theta}\sim\mathcal{U}([0,100,0,20,10,1,1,0,0,0,0.5,0,0,0,0],[80,600,4,60,60,7,12,2,3,0.7,2.5,4,3,70,300])$
Simulator:
The model is simulated using Gillespie’s stochastic simulation algorithm (SSA) over $200$ time steps. The key observables are the copy numbers of the complex $C$, activator $A$, and repressor $R$.
Dimensionality:
The parameter vector $\bm{\theta}\in\mathbb{R}^{15}$, and the output time series data has the shape $N\times S\times T$, where $N$ is the number of samples, $S=3$ is the number of observed species, and $T=200$ is the number of time steps. To reduce the dimensionality and make the data more manageable, we train a simple autoencoder. The autoencoder compresses the data from $N\times 3\times 200$ to $N\times 15$, effectively replacing the time series with a set of summary statistics.
Fixed Parameters:
The model is initialized with the following parameter values and species concentrations:
  • True parameters:
    $\alpha_{a}$ $=50,\quad\alpha_{a}^{\prime}=500,\quad\alpha_{r}=0.01,\quad\alpha_{r}^{\prime}=50,$
    $\beta_{a}$ $=50,\quad\beta_{r}=5,\quad\delta_{ma}=10,\quad\delta_{mr}=0.5,$
    $\delta_{a}$ $=1,\quad\delta_{r}=0.2,\quad\gamma_{a}=1,\quad\gamma_{r}=1,$
    $\gamma_{c}$ $=2,\quad\theta_{a}=50,\quad\theta_{r}=100.$
  • Initial Species Values:
    $D_{A}=1,\;D_{A}^{*}=0,\;M_{A}=0,\;D_{R}=1,\;D_{R}^{*}=0,\;M_{R}=0,\;C=10,\;A=10,\;R=10.$
[…]
Figure 7: Time series trajectories for three key species in the Vilar oscillator system: Complex (C), Activator (A), and Repressor (R). The darker line shows a reference trajectory obtained using the true parameter values, while the four lighter lines represent additional stochastic realizations under identical conditions. This visualization highlights the intrinsic variability of the system while preserving its consistent oscillatory behavior.
Appendix B Hyperparameters
In this section, we describe the hyperparameters and simulation environments employed across all methods. Training data for the SBI benchmark problems were generated using the sbibm [Lueckmann et al., 2021], except for the LV and SIR models, which were generated with the BayesFlow framework [Radev et al., 2023b]. For the HH model, we adopt the implementation provided by [Gloeckler et al., 2024], and for the Genetic Oscillator model, we follow the approach of [Akesson et al., 2022]. In this model, the simulated data is high dimensional, with a shape of $(N\times 3\times 200)$ representing three species observed over 200 timesteps. We compress this data using an autoencoder whose latent space is set to match the number of simulator parameters. For Simformer, we use hyperparameters published in their original works [Gloeckler et al., 2024]. For GATSBI, task-specific hyperparameters were taken from [Ramesh et al., 2022] and adjusted as needed based on each problem’s structure and computational demands.
For NPE, we follow [Lueckmann et al., 2021], which evaluated NPE across ten benchmark problems. As reported in Appendix H of sbibm paper, Neural Spline Flows (NSF) generally outperform Masked Autoregressive Flows (MAF), originally recommended for NPE. Consequently, we use the NSF implementation from the sbi package [Tejero-Cantero et al., 2020] as the density estimator in all experiments. The implementation of Automatic Posterior Transformation (APT, or SNPE-C) is also taken from the sbi package; as a sequential method, the simulation budget per round is computed by dividing the total budget by the number of rounds. For time series problems, a summary network is recommended for NPE/APT [Zeghal et al., 2022]. Accordingly, we modify and optimize the standard CNNEmbedding from the sbi package to convert the data into summary statistics with dimensions matching the parameters (4 for the LV test problem and 2 for the SIR task).
For the ConDiSim model, we optimized the hyperparameters for each of the ten benchmark problems from the sbibm library, as detailed in Table 3. Common hyperparameters were applied uniformly across all tasks to ensure training consistency and stability. We used the AdamW optimizer with a learning rate of $0.001$ and a weight decay of $0.0001$, along with a ReduceLROnPlateau scheduler to decrease the learning rate by a factor of $0.5$ if validation loss did not improve over five consecutive epochs. The noise schedule followed a cosine/quadratic scheme, with $\beta_{\text{min}}$ as $0.0001$ and $\beta_{\text{max}}$ as $0.02$, to regulate progressive noise addition during diffusion. Data standardization was conducted using StandardScaler, and an $80/20$ split was used for training and validation. Early stopping with a patience of $20$ epochs was applied to prevent overfitting. SiLU (Sigmoid Linear Unit) activations were used across all layers to enhance pattern capture and improve stability. This consistent training strategy enabled a systematic comparison of model performance across various batch sizes, diffusion steps, and hidden unit configurations.
Table 3: Hyperparameters for ConDiSim Model
Benchmark Batch Size Diffusion Steps Hidden Units SLCP 32 200 Linear, (512, 512, 512) SLCP with Distractors 32 1000 Linear, (512, 512, 512) Two Moons 32 50 Linear, (256, 256, 256) Gaussian Mixture 32 50 Linear, (256, 256, 256) Gaussian Linear Uniform 32 50 Linear, (256, 256, 256) SIR 32 50 Linear, (256, 256, 256) Bernoulli GLM 128 500 Linear, (256, 256, 256) Bernoulli GLM Raw 128 500 Linear, (512, 512, 512) Lotka-Volterra 128 200 Conv1D (32, 64, 128), Linear, (128, 128) HH Model 32 300 Linear, (256, 256, 256) Genetic Oscillator 64 400 Linear, (256, 256, 256, 256)
Appendix C Additional Results
We present plots of estimated posterior distributions for the remaining benchmark test problems below:
C.1 Posterior Evolution for Two Moons
The evolution of the approximated posterior for the Two Moons test problem is shown in Fig. 8.
Figure 8: Evolution of the posterior distribution for the Two Moons task, showing the transition from a broad, noisy distribution to a well-defined bi-modal structure as $T\to 0$.
(a) Hodgkin–Huxley: Approximated Posterior
(b) Posterior Predictive Voltage Traces
Figure 9: Simformer:Posterior Distributions for Hodgkin–Huxley model
C.2 Performance of Simformer on HH Model
Unlike the standard SBI benchmarks, where true posteriors are known for a known set of true parameters, the Hodgkin-Huxley (HH) model lacks an analytical posterior, preventing direct comparison via metrics like C2ST or MMD. For fair evaluation, we first benchmark CondiSim on 10 standard SBI tasks (refer Section 4.1) before demonstrating its real-world applicability to the HH model. Given a fixed parameter configuration $\theta^{*}=(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606)$, we infer posteriors using both CondiSim and Simformer (trained with 30k simulations). Posterior samples and posterior predictive checks for Simformer (Figure 9) and ConDiSim (Figure 4) confirm that both methods produce plausible voltage traces, with ConDiSim yielding tighter posterior support on several parameters.
Appendix D Posterior Plots for SBI benchmarking problems
In this section, we summarize the unique challenges posed by remaining SBI benchmark problem and highlight the performance of our model. Detailed posterior distributions are provided in the corresponding figures for visual reference.
In the Gaussian Mixture (GMM) problem, the main challenge is identifying the shared mean between two Gaussian distributions with different variances. Figure 10(a) demonstrates that our model effectively captures the primary modes and shapes of these distributions.
For Gaussian Linear and Gaussian Linear Uniform, the challenge is accurately modeling the posterior under both conjugate and non-conjugate settings. The Gaussian Linear Uniform variant complicates posterior inference with a uniform prior. Figure 11(a) and Figure 11(b) shows that our model effectively handles these differences.
In Bernoulli GLM and Bernoulli GLM Raw, the challenge involves estimating posteriors from Bernoulli observations with Gaussian priors. The "Raw" variant introduces complexity by using raw data instead of sufficient statistics. Figures 12(a) and 12(b) indicate accurate posterior modeling for both cases.
The SIR model aims to infer contact and recovery rates from noisy infection data over time. Figure 10(b) illustrates that our model robustly estimates the parameters despite the high variability in data.
Lastly, the Lotka-Volterra model is challenging due to its nonlinear dynamics and potential chaotic behavior. Figure 13 shows that our model successfully infers the underlying parameters, effectively navigating the model’s complexity.
These experiments demonstrate that ConDiSim effectively infers posterior distributions across all benchmark problems with high accuracy.
(a) GMM
(b) SIR Model
Figure 10: Posterior Plots: Comparison of generated vs. reference posterior distributions for the GMM and SIR models.
(a) Gaussian Linear Model
(b) Gaussian Linear Uniform Model
Figure 11: Posterior Distributions for Gaussian Linear and Gaussian Linear Uniform Models: Generated vs. Reference Posterior
(a) Bernoulli GLM
(b) Bernoulli GLM Raw
Figure 12: Posterior Distributions for Bernoulli GLM and Bernoulli GLM Raw: Generated vs. Reference Posterior
Figure 13: Lotka–Volterra Posterior: Visualization of the posterior for the Lotka–Volterra model. The true posterior mean is indicated by a vertical line (due to low variance).
Appendix E Simulation Based Calibration
Simulation-Based Calibration (SBC) is an essential metric for generative models in Simulation-Based Inference (SBI), as it verifies alignment between inferred posterior distributions and true parameter values, assisting in detecting potential biases in model-generated samples. For SBC, we generate $M=500$ prior samples, simulate corresponding observations, and draw $L=250$ posterior samples for each observation using the learned ConDiSim model. Calibration is assessed through rank statistics of true parameters within their posterior distributions, with empirical cumulative distribution functions (ECDFs) providing a robust, bin-free visualization of these ranks. Following [Säilynoja et al., 2022], ECDF-based SBC transforms posterior draws into fractional rank statistics, avoiding arbitrary binning and facilitating interpretation. A uniform ECDF indicates well-calibrated posteriors, while deviations suggest miscalibration: curves above the diagonal indicate overconfidence, and those below indicate underconfidence. Using BayesFlow [Radev et al., 2023b], we generate ECDF plots with 90% confidence interval bands as calibration benchmarks. Solid blue lines represent rank ECDFs of posterior draws, with shaded grey areas denoting the 90% confidence intervals. The SBC results for SBI benchmarking problems are presented below:
(a) SLCP
(b) SLCP Distractors
(c) Gaussian Linear
(d) Gaussian Linear Uniform
(e) Bernoulli GLM
(f) Bernoulli GLM Raw
(g) SIR
(h) Gaussian Mixture
(i) Lotka Volterra
Figure 16: ECDF results for the 30,000 simulation budget and the first run for each problem.
Appendix F SBI Metrics
F.1 Limitations of the C2ST Metric
The Classifier Two-Sample Test (C2ST) evaluates the similarity between a generated posterior $P$ and the true posterior $Q$ by training a classifier to differentiate their samples. A score near 0.5 indicates that the two distributions are similar, while higher values indicate discrepancies. Although widely used in SBI literature for its straightforward interpretability—providing a clear score between 0.5 and 1—the metric can be overly sensitive, as a single poorly estimated dimension may distort the results. Furthermore, its reliability depends significantly on the design and training of the classifier.
F.2 MMD
To overcome these limitations, we also consider the Maximum Mean Discrepancy (MMD), which measures the distance between the mean embeddings of $P$ and $Q$ in a reproducing kernel Hilbert space (RKHS). MMD is defined as,
$\mathrm{MMD}^{2}(P,Q;\mathcal{H})=\left\|\mathbb{E}_{x\sim P}[\phi(x)]-\mathbb{E}_{y\sim Q}[\phi(y)]\right\|_{\mathcal{H}}^{2},$ (23)
where $\phi$ is a feature mapping into the RKHS $\mathcal{H}$. A low MMD value indicates that the generated posterior closely approximates the true posterior. Unlike C2ST, MMD provides a robust, holistic measure of similarity between distributions without being overly influenced by discrepancies in a single dimension. A low MMD value indicates that the distributions $P$ and $Q$ are similar, while a larger value suggests significant differences.
Table 4: MMD metric for different methods across simulation budgets (mean $\pm$ std. dev.).
(a) Two Moons, Bernoulli GLM, and SIR
Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 FMPE 0.0133 $\pm$ 0.0197 0.0039 $\pm$ 0.0060 0.019 $\pm$ 0.0014 0.1076 $\pm$ 0.0283 0.0826 $\pm$ 0.0138 0.0746 $\pm$ 0.0247 0.2058 $\pm$ 0.0024 0.3618 $\pm$ 0.0048 0.1819 $\pm$ 0.0044 ConDiSim 0.0105 $\pm$0.0041 0.0104 $\pm$0.0048 0.0074 $\pm$0.0002 0.3157 $\pm$0.1771 0.1471 $\pm$0.0428 0.1981 $\pm$0.1348 0.7560 $\pm$0.1803 0.5754 $\pm$0.2392 0.6909 $\pm$0.1521
(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra
Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 FMPE 0.1598 $\pm$ 0.0501 0.1403 $\pm$ 0.0471 0.0930 $\pm$ 0.0561 0.0859 $\pm$ 0.0167 0.0831 $\pm$ 0.0181 0.0781 $\pm$ 0.0102 1.0537 $\pm$ 0.6277 1.0858 $\pm$ 0.5639 1.1696 $\pm$ 0.4336 ConDiSim 1.0688 $\pm$0.3057 0.6939 $\pm$0.3204 0.4105 $\pm$0.3082 0.1448 $\pm$0.0368 0.1435 $\pm$0.0822 0.1658 $\pm$0.1407 3.4480 $\pm$0.8115 3.4678 $\pm$0.3850 3.7210 $\pm$0.5691
(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.2242 $\pm$0.0296 0.1149 $\pm$0.0549 0.1733 $\pm$0.1082 0.4999 $\pm$0.1595 0.6476 $\pm$0.3050 0.7182 $\pm$0.2839 0.9044 $\pm$0.2353 0.7712 $\pm$0.1483 1.1197 $\pm$0.4795 0.7197 $\pm$0.5598 0.3375 $\pm$0.1473 0.7948 $\pm$0.5992 NPE 0.0604 $\pm$0.0260 0.0418 $\pm$0.0211 0.0554 $\pm$0.0347 0.0105 $\pm$0.0036 0.0074 $\pm$0.0018 0.0082 $\pm$0.0011 0.0258 $\pm$0.0098 0.0117 $\pm$0.0041 0.0121 $\pm$0.0051 0.4021 $\pm$0.0551 0.1632 $\pm$0.0684 0.2333 $\pm$0.0935 APT 0.0137 $\pm$0.0031 0.0108 $\pm$0.0088 0.0110 $\pm$0.0034 0.0075 $\pm$0.0013 0.0070 $\pm$0.0018 0.0080 $\pm$0.0008 0.0166 $\pm$0.0087 0.0138 $\pm$0.0053 0.0113 $\pm$0.0039 0.1405 $\pm$0.0347 0.0487 $\pm$0.0296 0.0287 $\pm$0.0166 Simformer 0.0009 $\pm$0.0007 0.0009 $\pm$0.0006 0.0007 $\pm$0.0010 0.0092 $\pm$0.0116 0.0037 $\pm$0.0035 0.0037 $\pm$0.0028 0.0014 $\pm$0.0006 0.0015 $\pm$0.0011 0.0008 $\pm$0.0003 0.0323 $\pm$0.0203 0.0025 $\pm$0.0034 0.0021 $\pm$0.0030 FMPE 0.1893 $\pm$ 0.0985 0.0208 $\pm$ 0.0078 0.0013 $\pm$ 0.0011 0.0511 $\pm$ 0.0053 $0$$\pm$ 0.0041 0.0497 $\pm$ 0.0034 0.1425 $\pm$ 0.0219 0.1261 $\pm$ 0.0206 0.1236 $\pm$ 0.0247 0.1573 $\pm$ 0.0543 0.1489 $\pm$ 0.0528 0.1489 $\pm$ 0.0528 ConDiSim 0.1604 $\pm$0.0599 0.0720 $\pm$0.0124 0.0974 $\pm$0.0505 0.0471 $\pm$0.0126 0.0387 $\pm$0.0129 0.0290 $\pm$0.0038 0.0335$\pm$0.0139 0.0273 $\pm$0.0079 0.0285 $\pm$0.0098 0.2760 $\pm$0.0493 0.1954 $\pm$0.1029 0.1572 $\pm$0.0726

### v2

@@
 Abstract
-We present a conditional diffusion model - ConDiSim, for simulation-based inference of complex systems with intractable likelihoods.
+We present ConDiSim, a conditional diffusion model for simulation-based inference in complex systems with intractable likelihoods.
 ConDiSim leverages denoising diffusion probabilistic models to approximate posterior distributions, consisting of a forward process that adds Gaussian noise to parameters, and a reverse process learning to denoise, conditioned on observed data.
@@
 ConDiSim is evaluated across ten benchmark problems and two real-world test problems, where it demonstrates effective posterior approximation accuracy while maintaining computational efficiency and stability in model training.
-ConDiSim offers a robust and extensible framework for simulation-based inference, particularly suitable for parameter inference workflows requiring fast inference methods.
+ConDiSim provides a robust and extensible framework for simulation-based inference, well suited to parameter estimation tasks that demand fast methods for handling noisy, time series observations.
 1 Introduction
@@
 Calibrating these parameters to ensure that simulator outputs align with observed data constitutes an inverse problem, formally defined within the framework of simulation-based inference (SBI) [Cranmer et al., 2020].
-Solving this inverse problem involves addressing uncertainties arising from model stochasticity and potential multi-valuedness, where different sets of parameter values can produce similar observations or similar parameters may lead to varied outputs.
-Additionally, parameter inference becomes increasingly complex when simulators operate as ‘black boxes’ with intractable likelihood functions, rendering traditional likelihood-based Bayesian methods impractical [Sisson et al., 2018].
+Such inverse problems are often ill-posed, meaning they may lack unique or stable solutions, especially when the simulator is stochastic or when different parameter configurations produce similar observations [Sisson et al., 2018].
+To address these challenges, the Bayesian formulation of SBI places a prior distribution over the parameters and infers a posterior distribution conditioned on observed data.
+This posterior captures the full set of plausible solutions and enables principled uncertainty quantification, offering a more robust alternative to relying on single-point estimates.
 Recent advances in machine learning have introduced powerful generative models capable of approximating complex data distributions without explicit likelihood computations.
-Among these, diffusion models have emerged as a particularly promising class [Sohl-Dickstein et al., 2015].
-Originally developed for image and signal processing, diffusion models iteratively denoise data starting from pure noise, effectively learning data distributions through a sequence of transformations.
-Compared to alternative generative frameworks like Variational Autoencoders (VAEs) [Kingma and Welling, 2014], Generative Adversarial Networks (GANs) [Goodfellow et al., 2014], and flow-based models [Rezende and Mohamed, 2015], diffusion models offer significant advantages.
-They avoid issues like posterior collapse in VAEs and mode collapse in GANs, and are not constrained by invertible transformations as in flow-based models, enabling them to better model complex, high-dimensional, and multi-modal distributions.
-Diffusion models also demonstrate superior training stability and come with theoretical convergence guarantees due to their foundation in stochastic differential equations [Ho et al., 2020].
-Their flexible architectures and sampling processes allow practitioners to balance sample quality and generation speed, enhancing their applicability in various contexts, including SBI.
-Applying diffusion models to SBI leverages their ability to implicitly learn the conditional distribution of simulator parameters given observations, circumventing the need for explicit likelihood evaluations.
-This is particularly advantageous for inverse problems involving black-box simulators.
-Moreover, diffusion models naturally support amortized inference: a single, globally trained model approximates the posterior distribution across multiple observations, eliminating the need for separate optimization for each new observation [Kingma and Welling, 2014].
-This capability is especially valuable when rapid inference is required.
-However, integrating diffusion models into SBI presents challenges.
-Their iterative sampling process can be computationally intensive, and their performance is sensitive to hyperparameters like noise schedules and network architecture.
-Addressing these issues is crucial for enhancing the practical utility of diffusion-based SBI methods [Nichol and Dhariwal, 2021].
-Contribution: We present a conditional diffusion model (ConDiSim) for simulation-based inference that balances model expressiveness with training efficiency and simplicity.
-Experiments on benchmark tasks and real-world datasets show the effectiveness of the approach in recovering complex posterior distributions, while incurring minimal computational overhead.
-Further, simulation-based calibration checks validate the reliability and robustness of the proposed parameter inference approach.
-The paper is organized as follows.
+Among these, diffusion models have emerged as a particularly promising class [Sohl-Dickstein et al., 2015, Ho et al., 2020, Song et al., 2020].
+They model data by learning to reverse a fixed noising process, gradually denoising samples from pure noise.
+Moreover, diffusion models naturally support amortized inference [Kingma and Welling, 2014]: a single trained model approximates posteriors for any observation without per-instance optimization, enabling faster inference for new data.
+This work explores diffusion models for addressing common SBI challenges: limited simulation budgets from expensive simulators (particularly in systems biology and epidemiology with stiff ODE/PDE solvers), noisy and stochastic outputs, multimodal posteriors from nonlinearities and identifiability issues, and the need for architectures that integrate parameters, timesteps, and observations to recover diverse posterior structures.
+ConDiSim is demonstrated to be well-suited for SBI problems involving learning low-dimensional posteriors from noisy, high-dimensional discrete-time observations.
+Contributions: (i)We adapt conditional Denoising Diffusion Probabilistic Models (DDPMs) for SBI to learn complex posterior distributions under limited simulation budgets; (ii) ConDiSim introduces a novel reverse diffusion architecture for SBI that employs time-dependent Feature-wise Linear Modulation (FiLM) to jointly condition on noisy parameters, diffusion timesteps, and observations for posterior estimation; (iii) We empirically evaluate ConDiSim systematically across benchmarks and real-world SBI tasks, analyzing performance under limited simulation budgets, noisy outputs, and multimodal posteriors.
+Outline: The paper is organized as follows.
 Section 2 outlines the related work.
-In Section 3, we formally define the SBI problem, followed by a detailed explanation of the ConDiSim model in Section 3.1 and Section 3.2 details our amortized inference strategy.
-A comprehensive comparison on benchmark, and real-world test problems is presented in Section 4, with additional results provided in the Appendix.
-Section 5 discusses the results, while Section 6 summarizes the key findings and suggests future directions.
+Section 3 formally defines the SBI problem.
+Section 3.1 introduces the ConDiSim model, and Section 3.2 details its conditioning mechanism and sampling procedure, including the proposed architecture.
+Section 4 presents a comprehensive evaluation on benchmark and real-world problems, with additional results provided in the Appendix.
+Section 5 presents a discussion of the results, while Section 6 concludes the paper.
 2 Related Work
-Simulation-based inference (SBI) has advanced significantly in recent years, evolving from traditional likelihood-free methods to modern deep learning approaches.
-Classical techniques like Approximate Bayesian Computation (ABC) [Beaumont et al., 2002] and its sequential variants such as Sequential Monte Carlo-ABC (SMC-ABC) [Toni et al., 2009], provide mathematical convergence guarantees but often struggle with high-dimensional problems [Sisson et al., 2018].
-Although neural network-learned summary statistics [Jiang et al., 2017, Wrede et al., 2022] have contributed to mitigating this limitation, they introduce additional computational complexity.
-Early advancements in neural network-based amortized inference for SBI were pioneered by Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016], which introduced conditional normalizing flows to model the posterior distribution, significantly improving accuracy and scalability over traditional methods.
-Sequential extensions like SNPE-A [Papamakarios and Murray, 2018], SNPE-B [Lueckmann et al., 2017], and SNPE-C (also known as Automatic Posterior Transformation or APT) [Greenberg et al., 2019] enhanced efficiency by iteratively refining the model to focus on relevant parameter regions.
-However, these sequential methods require multiple training rounds and are constrained by the invertible transformations required in flow-based models.
-Alternative sequential algorithms have also been proposed, such as Sequential Neural Ratio Estimation (SNRE) [Hermans et al., 2020], which learns the likelihood-to-evidence ratio, and Sequential Neural Likelihood Estimation (SNLE) [Papamakarios et al., 2019], which approximates the likelihood function directly.
-Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] leverages adversarial networks to model complex distributions without explicit likelihood calculations.
-While effective for challenging inference problems, GATSBI faces issues common to adversarial training, such as instability and mode collapse [Arjovsky and Bottou, 2017, Arjovsky et al., 2017], which can limit robustness and complicate training, and it struggles in low-dimensional spaces as noted in [Ramesh et al., 2022].
-Recently, Simformer [Gloeckler et al., 2024] combined score-based diffusion models with transformer architectures to accurately learn complex, high-dimensional posteriors, offering high flexibility albeit with substantial computational overhead.
-In contrast, our proposed method, ConDiSim, utilizes denoising-based diffusion models with a simpler architecture, maintaining competitive performance while being computationally efficient.
-This makes it suitable for applications such as model exploration, that may involve numerous parameter inference problems with inference time being a key consideration [Wrede and Hellander, 2019].
-In this study, we evaluate CondiSim against four distinct simulation-based inference (SBI) approaches: GATSBI [Ramesh et al., 2022], APT [Greenberg et al., 2019], NPE [Papamakarios and Murray, 2016], and Simformer [Gloeckler et al., 2024].
-These methods represent a diverse spectrum of generative modeling techniques—namely, flow-based, adversarial, diffusion, one-shot, and sequential frameworks—adapted for SBI.
-This comprehensive selection facilitates an evaluation that emphasizes efficiency and scalability rather than focusing solely on the accuracy of posterior approximations.
-For further details on deep learning methodologies applied to SBI, readers are referred to [Zammit-Mangion et al., 2024, Radev et al., 2023a].
+Simulation-based inference has advanced significantly in recent years, evolving from traditional statistical likelihood-free methods to contemporary deep learning frameworks.
+Classical techniques like Approximate Bayesian Computation (ABC) [Beaumont et al., 2002] and Sequential Monte Carlo-ABC (SMC-ABC) [Toni et al., 2009] offer theoretical convergence guarantees but may scale poorly to high dimensional problems [Sisson et al., 2018].
+While neural summary statistics [Jiang et al., 2017, Wrede et al., 2022] can reduce dimensionality, they introduce trade-offs: additional computational burden and the possibility of losing information necessary for accurate posterior estimation.
+Neural Posterior Estimation (NPE) [Papamakarios and Murray, 2016] pioneered the use of deep generative models for amortized SBI, introducing conditional normalizing flows to model the posterior distribution, substantially improving accuracy and scalability over traditional methods.
+Sequential extensions like SNPE-A [Papamakarios and Murray, 2018], SNPE-B [Lueckmann et al., 2017], and SNPE-C (also known as Automatic Posterior Transformation or APT) [Greenberg et al., 2019] enhanced accuracy by iteratively refining the model to focus on relevant parameter regions.
+However, these sequential methods require multiple rounds of training and are constrained by the invertible transformations required in flow-based models.
+Recently, Flow Matching Posterior Estimation (FMPE) [Wildberger et al., 2023] introduced continuous normalizing flows based on conditional vector fields, overcoming architectural constraints and efficiently scaling to high-dimensional inference tasks.
+However, FMPE requires ODE solvers at inference time, introducing additional computational costs, and its performance depends critically on the choice of flow interpolant, which can be challenging to optimize under tight simulation budgets.
+Beyond flow-based methods, other deep generative approaches to SBI include adversarial models.
+Generative Adversarial Training for SBI (GATSBI) [Ramesh et al., 2022] uses adversarial networks to model posteriors but suffers from training instability, mode collapse [Arjovsky and Bottou, 2017, Arjovsky et al., 2017], and poor performance in low-dimensional settings [Ramesh et al., 2022].
+More recently, Simformer [Gloeckler et al., 2024] combined score-based diffusion models with a transformer architecture, enabling high-fidelity posterior estimation but at the expense of increased computational cost.
+Score-based diffusion models have also been explored in the tall-data setting of SBI, where multiple observations are combined efficiently using amortized score networks [Linhart et al., 2024].
+In contrast, ConDiSim uses DDPMs with a lightweight, FiLM-based architecture that avoids heavy attention mechanisms while achieving competitive performance with significantly lower inference and training overhead.
+This makes it especially suitable for settings like iterative model exploration, where many inference queries must be resolved quickly [Wrede and Hellander, 2019].
+In this study, we evaluate ConDiSim against five distinct SBI approaches: NPE [Papamakarios and Murray, 2016], APT [Greenberg et al., 2019], GATSBI [Ramesh et al., 2022], FMPE [Wildberger et al., 2023], and Simformer [Gloeckler et al., 2024].
+Together, these methods span a wide spectrum of generative modeling paradigms, including discrete flow-based methods in both one-shot and sequential forms, continuous flow-based approaches, adversarial models, and diffusion models in both discrete and continuous formulations.
+This selection enables a balanced evaluation that emphasizes efficiency and scalability in addition to posterior accuracy.
+For a broader overview of deep learning approaches to SBI, see [Zammit-Mangion et al., 2024, Radev et al., 2023a].
 3 Simulation-Based Inference
@@
 $p(\bm{\theta}\mid\mathbf{y})\propto p(\mathbf{y}\mid\bm{\theta})\,p(\bm{\theta}),$ (1)
-where $p(\mathbf{y}\mid\bm{\theta})$ is the likelihood function and $p(\bm{\theta})$ is the prior distribution.
-In many complex models, especially those involving stochastic simulators, the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is intractable or computationally expensive to evaluate, rendering traditional inference methods like Markov Chain Monte Carlo (MCMC) algorithms [Andrieu et al., 2003] infeasible due to their reliance on explicit likelihood evaluations.
-When the likelihood function is inaccessible, a popular strategy is to generate empirical samples using a stochastic simulator $\mathcal{M}(\bm{\theta},\bm{\xi})$ that produces synthetic data $\mathbf{y}$ given parameters $\bm{\theta}$ and random variables $\bm{\xi}$ representing system noise.
-Sampling $\bm{\theta}$ from a prior distribution $p(\bm{\theta})$ informed by domain knowledge or chosen to be noninformative and generating the corresponding $\mathbf{y}$ through the simulator, we obtain samples from the joint distribution $p(\bm{\theta},\mathbf{y})$.
-These joint samples allow us to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$ using generative modeling techniques that learn the inverse mapping from observations $\mathbf{y}$ to parameters $\bm{\theta}$ without requiring explicit likelihood calculations.
-3.1 ConDiSim: A Conditional Diffusion Framework for Simulation-Based Inference
-We now introduce ConDiSim — a denoising conditional diffusion model [Ho et al., 2020] that systematically addresses the challenge of intractable likelihoods by leveraging samples from the joint distribution $p(\bm{\theta},\mathbf{y})$ , enabling efficient and accurate posterior inference without the need for explicit likelihood evaluations.
-The model consists of a forward diffusion process that gradually adds Gaussian noise to the parameters $\bm{\theta}$, transforming them into a tractable noise distribution.
-The reverse process then denoises these parameters, conditioned on observed data $\mathbf{y}$, to approximate the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$.
-Figure 1 provides a schematic overview of the ConDiSim model.
+where $p(\mathbf{y}\mid\bm{\theta})$ is the likelihood and $p(\bm{\theta})$ is the prior distribution.
+In many scientific models, particularly those involving stochastic simulators, the likelihood function $p(\mathbf{y}\mid\bm{\theta})$ is either intractable or computationally expensive to evaluate, rendering likelihood-based methods such as Markov Chain Monte Carlo (MCMC) infeasible [Andrieu et al., 2003].
+When the likelihood is inaccessible, a common strategy is to sample parameters from a prior, $\bm{\theta}\sim p(\bm{\theta})$, where $p(\bm{\theta})$ may be informed by domain knowledge or chosen to be noninformative, such as a broad uniform prior.
+For each sampled parameter, a corresponding observation is generated via the simulator, $\mathbf{y}\sim p(\mathbf{y}\mid\bm{\theta})$.
+This yields joint samples $(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})$, which can then be used to train generative models that approximate the posterior $p(\bm{\theta}\mid\mathbf{y})$ without requiring explicit likelihood evaluation.
+3.1 ConDiSim: Conditional Diffusion Framework for SBI
+We adapt conditional denoising diffusion probabilistic models [Ho et al., 2020] to SBI, where the goal is to approximate the posterior $p(\bm{\theta}\mid\mathbf{y})$ in likelihood-free settings using joint samples $(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})$.
+The model employs a forward diffusion process that gradually corrupts the parameters $\bm{\theta}$ with Gaussian noise until they follow a tractable prior distribution.
+The reverse process then learns to denoise these corrupted parameters, conditioned on observations $\mathbf{y}$, thereby recovering samples from the posterior distribution $p(\bm{\theta}\mid\mathbf{y})$.
 In the forward diffusion process, we construct a discrete-time Markov chain $\{\bm{\theta}_{t}\}_{t=0}^{T}$, initialized at $\bm{\theta}_{0}=\bm{\theta}$.
-At each step $t$, the state $\bm{\theta}_{t}$ is derived from $\bm{\theta}_{t-1}$ by perturbing it with Gaussian noise according to the transition kernel:
+At each step $t$, the new state $\bm{\theta}_{t}$ is derived from the old state $\bm{\theta}_{t-1}$ by perturbing it with Gaussian noise according to the transition kernel:
 $q(\bm{\theta}_{t}\mid\bm{\theta}_{t-1})=\mathcal{N}\big(\bm{\theta}_{t};\,\sqrt{1-\beta_{t}}\,\bm{\theta}_{t-1},\,\beta_{t}\mathbf{I}\big),$ (2)
@@
 The joint reverse distribution is formulated as:
-$p(\bm{\theta}_{0:T}\mid\mathbf{y})=p(\bm{\theta}_{T})\,\prod_{t=1}^{T}p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}),$ (5)
-Training the reverse model involves maximizing the log-posterior of the latent variables given the observations:
+$p(\bm{\theta}_{0:T}\mid\mathbf{y})=p(\bm{\theta}_{T})\,\prod_{t=1}^{T}p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}).$ (5)
+Training the reverse model involves maximizing the log‑posterior of the latent variables, given the observations:
 $\mathcal{L}_{\mathrm{ConDiSim}}=\max\log p(\bm{\theta}_{0}\mid\mathbf{y}),$ (6)
-[…]
-However, evaluating $\log p(\bm{\theta}_{0}\mid\mathbf{y})$ directly is intractable, as it requires integrating over the full trajectory of intermediate latent states.
-To address this, we employ a variational approach by introducing the joint forward distribution $q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})$.
-Using Jensen’s inequality, the lower bound on the log-posterior is given by:
+$\text{where,}\quad\log p(\bm{\theta}_{0}\mid\mathbf{y})=\log\int p(\bm{\theta}_{0:T}\mid\mathbf{y})\,d\bm{\theta}_{1:T}.$ (7)
+However, evaluating $\log p(\bm{\theta}_{0}\mid\mathbf{y})$ directly is intractable, as it entails integrating over the full trajectory of intermediate latent states.
+Instead, we employ a variational approach by introducing the joint forward distribution $q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})$.
+Using Jensen’s inequality, the lower bound on the log-posterior is:
 $\log p(\bm{\theta}_{0}\mid\mathbf{y})\;\geq\;\mathbb{E}_{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\!\biggl[\log\frac{p(\bm{\theta}_{0:T}\mid\mathbf{y})}{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\biggr].$ (8)
 Applying Eqs. (4) and (5) to Eq. (8) leads to the following decomposition:
-$\mathcal{L}_{\mathrm{ConDiSim}}=\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{prior}}+\mathcal{L}_{\mathrm{denoise}}$ (9)
+$\mathcal{L}_{\mathrm{ConDiSim}}=\mathcal{L}_{\mathrm{recon}}+\mathcal{L}_{\mathrm{prior}}+\mathcal{L}_{\mathrm{denoise}},$ (9)
 where, the reconstruction term, $\mathcal{L}_{\mathrm{recon}}=\mathbb{E}_{q}[\log p(\bm{\theta}_{0}\mid\bm{\theta}_{1},\mathbf{y})]$, evaluates how well the original parameters are recovered from the final denoised state.
@@
 The denoising term, $\mathcal{L}_{\mathrm{denoise}}=-\sum_{t=2}^{T}\mathbb{E}_{q}[D_{\mathrm{KL}}(q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\,\|\,p(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y}))]$, aligns reverse transitions with the forward process.
-For sufficiently large $T$, the reconstruction and prior terms vanish, leaving the denoising KL as the dominant training objective:
+For sufficiently large $T$, the reconstruction and prior terms vanish [Ho et al., 2020], leaving the denoising KL as the dominant training objective (i.e., $\mathcal{L}\approx\mathcal{L}_{\text{denoise}}$):
 $\mathcal{L}=\sum_{t=2}^{T}\mathbb{E}_{q}\bigl[D_{\mathrm{KL}}\bigl(q(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\bm{\theta}_{0})\;\|\;p_{\phi}(\bm{\theta}_{t-1}\mid\bm{\theta}_{t},\mathbf{y})\bigr)\bigr].$ (10)
@@
 To eliminate the explicit dependence on $\bm{\theta}_{0}$, we substitute $\bm{\theta}_{0}=(\bm{\theta}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\bm{\epsilon})/\sqrt{\bar{\alpha}_{t}}$ (with $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$) into Eq.
-(3), thereby reparameterizing $\bm{\mu}_{q}$ to yield a noise-parameterized expression for the mean:
+(3), reparameterizing $\bm{\mu}_{q}$ to yield a noise-parameterized expression for the mean:
 $\bm{\mu}_{q}(\bm{\theta}_{t},t)$ $=\tfrac{1}{\sqrt{\alpha_{t}}}\Bigl(\bm{\theta}_{t}-\tfrac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}\Bigr).$ (12)
-Assuming the learned reverse process preserves the covariance structure ($\bm{\Sigma}_{p}=\bm{\Sigma}_{q}$), the KL divergence in Eq. (10) reduces to the squared Euclidean distance between the means:
+Assuming the learned reverse process preserves the covariance structure ($\bm{\Sigma}_{p}=\bm{\Sigma}_{q}$) [Ho et al., 2020], the KL divergence in Eq.
+(10) reduces to the squared Euclidean distance between the means:
 $\mathcal{L}=\sum_{t=2}^{T}\mathbb{E}_{q(\bm{\theta}_{1:T}\mid\bm{\theta}_{0})}\Bigl[\big\lVert\bm{\mu}_{q}(\bm{\theta}_{t},\bm{\theta}_{0})\;-\;\bm{\mu}_{p}(\bm{\theta}_{t},\mathbf{y},t)\big\rVert^{2}\Bigr].$ (13)
@@
 $\bm{\mu}_{p}(\bm{\theta}_{t},\mathbf{y},t)$ $=\tfrac{1}{\sqrt{\alpha_{t}}}\Bigl(\bm{\theta}_{t}-\tfrac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\Bigr),$ (14)
-where, $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ is neural network that, given the noisy latent $\bm{\theta}_{t}$, conditioning $\mathbf{y}$, and time step $t$, predicts the noise.
-Substituting the mean parameterizations from Eqs.(12) and (14) into the loss of Eq.(13) and discarding all constant factors, yields the compact noise–prediction objective:
-$\mathcal{L}_{\text{ConDiSim}}=\mathbb{E}_{\,t,\,\bm{\epsilon},\,(\bm{\theta}_{0},\mathbf{y})\sim p(\theta,y)}\Bigl{[}\big{\lVert}\bm{\epsilon}-\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\big{\rVert}^{2}\Bigr{]}.$ (15)
-Figure 1: The ConDiSim Model: (a) The forward diffusion process incrementally adds Gaussian noise to the parameter vector ($\bm{\theta}$), generating a sequence of latent states ($\{\bm{\theta}_{t}\}$); (b) The reverse diffusion process reconstructs the posterior distribution ($p(\bm{\theta}\mid\mathbf{y})$) by leveraging the observed data ($\mathbf{y}$) for denoising.
-3.2 Amortized Inference with ConDiSim
-In our framework, we implement the reverse diffusion process as an instance of amortized inference.
-For a given observation $\mathbf{y_{o}}$, the reverse process is initialized by drawing $N$ latent variables $\bm{\theta}_{T}^{(i)}$ from the standard normal distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ for $i=1,\dots,N$.
-These latent samples are then conditioned on $\mathbf{y_{o}}$ and iteratively refined through the learned reverse diffusion process.
-Specifically, at each reverse time step, the transition is characterized by,
-$\bm{\theta}_{t-1}^{(i)}\mid\bm{\theta}_{t}^{(i)},\bm{y_{o}}\sim\mathcal{N}\!\left(\bm{\mu}_{\phi}\!\left(\bm{\theta}_{t}^{(i)},\mathbf{y_{o}},t\right),\,\bm{\Sigma}_{t}\right),$
-where $\bm{\mu}_{\phi}$ and $\bm{\Sigma}_{t}$ represent the learned mean and covariance parameters that determine the probabilistic mapping between consecutive latent states in the reverse process.
-After iterating from $t=T$ down to $t=0$, the process yields refined latent samples $\bm{\theta}_{0}^{(i)}$ that approximate the target posterior distribution $p(\bm{\theta}\mid\mathbf{y_{o}})$.
-This method efficiently produces high-quality posterior samples suitable for applications such as parameter estimation and uncertainty quantification.
+where, $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ is a neural network that, given the noisy latent $\bm{\theta}_{t}$, conditioning $\mathbf{y}$, and time step $t$, predicts the noise.
+Substituting the mean parameterizations from Eqs.(12) and (14) into the loss of Eq.(13) and discarding all constant factors, yields the compact noise prediction objective:
+$\mathcal{L}_{\text{ConDiSim}}\approx\mathbb{E}_{\,t,\,\bm{\epsilon},\,(\bm{\theta},\mathbf{y})\sim p(\bm{\theta},\mathbf{y})}\Bigl[\big\lVert\bm{\epsilon}-\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)\big\rVert^{2}\Bigr].$ (15)
+Figure 1: ConDiSim Architecture: Each diffusion block is FiLM-modulated by timestep and observation embeddings to estimate the noise injected at step $t$ during forward diffusion.
+3.2 Conditioning and Sampling
+The objective in Eq.
+(15) requires a denoising network that predicts the noise $\bm{\epsilon}$ from the noisy state $\bm{\theta}_{t}$, conditioned on the observation $\mathbf{y}$ and the diffusion timestep $t$.
+Preserving this dual conditioning throughout the denoising trajectory is essential to align with $p(\bm{\theta}\mid\mathbf{y})$.
+A naive scheme that feeds $\bm{\theta}_{t};\,\mathbf{y};\,t]$ into a standard network often performs poorly in SBI, where $\dim(\mathbf{y})\!\gg\!\dim(\bm{\theta})$, input-level concatenation forces the network to disentangle $\mathbf{y}$ and $t$ from $\bm{\theta}_{t}$ in a single step, yielding an ill-conditioned coupling that dilutes the posterior signal and destabilizes training.
+To address this, the proposed ConDiSim (Fig.
+1) model takes $(\bm{\theta}_{t},t,\mathbf{y})$ as inputs; while a condition encoder $f_{y}(\mathbf{y})$ and a timestep encoder $f_{t}(t)$ produce embeddings that are concatenated into a context vector $\mathbf{c}=[\,f_{y}(\mathbf{y});\;f_{t}(t)\,]$.
+A lightweight MLP maps $\mathbf{c}$ to per-block FiLM parameters $(\bm{\gamma}_{\ell}(\mathbf{c}),\bm{\beta}_{\ell}(\mathbf{c}))$, which modulate each diffusion block via $\operatorname{FiLM}_{\ell}(\mathbf{h}_{\ell})=\bm{\gamma}_{\ell}(\mathbf{c})\odot\mathbf{h}_{\ell}+\bm{\beta}_{\ell}(\mathbf{c})$ [Perez et al., 2018].
+This ensures that hidden features and thus the noise predictor $\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)$ remain observation-aware and time-aware at every step, avoiding the signal dilution of naive concatenation.
+Given this architecture, to perform amortized inference we draw $N$ initial states $\bm{\theta}_{T}^{(i)}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and evolve them backward for $t=T,\ldots,1$ :
+$\bm{\theta}_{t-1}^{(i)}=\frac{1}{\sqrt{\alpha_{t}}}\!\left(\bm{\theta}_{t}^{(i)}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t}^{(i)},\mathbf{y},t)\right)+\sigma_{t}\,\mathbf{z}_{t},$ (16)
+where $\mathbf{z}_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}=\sqrt{\beta_{t}}$.
+To balance fidelity to $\mathbf{y}$ with sample diversity, we apply classifier-free guidance by interpolating conditional and unconditional predictions, $\hat{\bm{\epsilon}}_{\phi}=(1+\lambda)\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},\mathbf{y},t)-\lambda\,\bm{\epsilon}_{\phi}(\bm{\theta}_{t},t),$ with guidance scale $\lambda\geq 0$.
+The final set $\{\bm{\theta}_{0}^{(i)}\}_{i=1}^{N}$ provides approximate samples from $p(\bm{\theta}\mid\mathbf{y})$ [Ho and Salimans, 2021].
 4 Experiments
-We evaluate ConDiSim on 10 benchmark problems from (sbibm) suite [Lueckmann et al., 2021] and two real-world problems.
-For each benchmark, we train with simulation budgets of 10 0001000010\,00010 000, 20 0002000020\,00020 000, and 30 0003000030\,00030 000 simulations, and infer the posterior by drawing 10 0001000010\,00010 000 samples from the model.
-Posterior quality is assessed with Maximum Mean Discrepancy (MMD), the Classifier Two-Sample Test (C2ST), and Empirical Cumulative Distribution Function (ECDF) diagnostics implemented in BayesFlow [Radev et al., 2023b].
-ECDF comparisons were performed using the BayesFlow framework [Radev et al., 2023b]; further details and ECDF plots for various benchmarks are provided in Appendix E.
-Complete hyperparameter configurations for all methods are detailed in Appendix B.
+We evaluate ConDiSim on 10 benchmark problems from the sbibm suite [Lueckmann et al., 2021], and two real‑world problems.
+For each benchmark, we constrain the training to simulation budgets of 10,00010{,}000, 20,00020{,}000, and 30,00030{,}000 simulations, and infer the posterior by drawing 10,00010{,}000 samples from the learned model.
+We assess posterior quality using Maximum Mean Discrepancy (MMD), the Classifier Two-Sample Test (C2ST) [Friedman, 2003], and Empirical Cumulative Distribution Function (ECDF) diagnostics from the BayesFlow framework [Radev et al., 2023b].
+C2ST results are reported in the main text; MMD scores and full ECDF plots for all benchmarks are provided in Appendix F.
+Complete hyperparameter configurations for all methods are detailed in Appendix C.
 (a) Posterior Plot
@@
 Figure 2: Two Moons: Comparison of generated and reference posterior distributions, with the ECDF plot representing the cumulative distribution of fractional rank statistics from posterior draws.
-(a) SLCP
-(b) SLCP Distractors
-Figure 3: SLCP: Comparison of generated vs. reference posteriors for SLCP and SLCP Distractors
+(a) SLCP.
+(b) SLCP Distractors.
+Figure 3: SLCP: Comparison of generated vs. reference posteriors for SLCP and SLCP Distractors.
+(a) Hodgkin–Huxley: Approximated Posterior
+(b) Posterior Predictive Voltage Traces
+Figure 4: Posterior Distributions for Hodgkin–Huxley model.
+(a) Genetic Oscillator: Approximated Posterior
+(b) Posterior Predictive Trajectories.
+Figure 5: Posterior Distributions for Genetic Oscillator model.
 4.1 Benchmark Posterior Analysis
-In this section, we analyze three representative benchmarks—Two Moons, SLCP, and SLCP Distractors—selected for their distinct challenges in posterior inference.
+In this section, we analyze three representative benchmarks: Two Moons, SLCP, and SLCP Distractors, selected for their distinct and complementary challenges in posterior inference.
 The Two Moons benchmark poses the challenge of capturing a bi-modal posterior with a distinctive crescent geometry.
@@
 The ECDF plots for $\theta_{1}$ and $\theta_{2}$ (Figure 2(b)) demonstrate that the inferred distributions are well-calibrated, with the majority of rank statistics falling within the 90% confidence bands; a minor deviation for $\theta_{1}$ suggests a slight degree of overconfidence.
-The SLCP benchmark encompasses a range of parameter behaviors - including uniform, Gaussian, multi-modal, and skewed distributions.
+The SLCP benchmark encompasses a range of parameter behaviors including uniform, Gaussian, multi-modal, and skewed distributions.
 Its variant, SLCP Distractors, further tests the model’s robustness by incorporating uninformative noise dimensions, thereby simulating scenarios with heavily corrupted data.
@@
 The rank ECDF plots for both SLCP variants (Figures 14(a) and 14(b) in the Appendix) also remain centered within the confidence intervals, confirming that the model reliably captures parameter variability even under noisy conditions.
-Detailed descriptions of these benchmarks, are provided in Appendix A.
-Table 1 presents the C2ST scores for the 10 benchmarking problems.
-For the Two Moons task, ConDiSim achieves C2ST performance comparable to Simformer, APT, and NPE, while requiring the shortest training time among the evaluated models (see Table 2).
-Simformer and APT perform well on most test problems – Simformer leveraging its transformer architecture with cross-attention, while APT is able to iteratively refine the posterior approximation.
-The SLCP problem remains particularly challenging, with all methods struggling to accurately infer the true posterior.
-It is important to note that the C2ST metric is sensitive to parameter estimation errors and can be difficult to interpret in high-dimensional settings, as a single poorly estimated dimension may disproportionately affect the overall score.
-For instance, as illustrated in Figure 3(a), ConDiSim encounters difficulties with the final dimension, despite accurately capturing the other dimensions, which is reflected in an overall C2ST score of 0.82.
+Detailed descriptions of these benchmarks, are provided in Appendix B.
+Table 1 reports C2ST scores across the 10 benchmarks.
+On tasks such as Two Moons, ConDiSim achieves performance comparable to Simformer, APT, and NPE, while requiring substantially shorter training times (Table 2).
+Although ConDiSim does not always attain the top C2ST score, the method offers a favorable trade-off between efficiency and posterior quality, which is crucial under limited simulation budgets.
+Notably, the SLCP task remains challenging for all methods, highlighting the inherent difficulty of accurately recovering high-dimensional, multi-modal posteriors.
+Moreover, C2ST can be overly sensitive to small errors in individual dimensions, making it less reliable in high-dimensional settings.
+For example, as shown in Figure 3(b), ConDiSim closely matches the true posterior in most dimensions, with deviations concentrated in the final parameter.
 To provide additional insight, we report the MMD metric for each benchmark problem in the Supplementary Table 4.
-(a) Hodgkin–Huxley: Approximated Posterior
-(b) Posterior Predictive Voltage Traces
-Figure 4: Posterior Distributions for Hodgkin–Huxley model
-(a) Genetic Oscillator: Approximated Posterior
-(b) Posterior Predictive Trajectories
-Figure 5: Posterior Distributions for Genetic Oscillator model
 4.2 Application to Real-World Scenarios
@@
 First, the Hodgkin-Huxley (HH) model [Gloeckler et al., 2024] which simulates neuronal action potentials by modeling the membrane voltage dynamics governed by sodium, potassium, and leak currents.
-Second, we consider the Genetic Oscillator model which is a high-dimensional gene regulatory network that exhibits oscillatory behavior, driven by a positive-negative feedback loop mimicking a circadian clock [Vilar et al., 2002].
+Second, we consider a Genetic Oscillator model which is a relatively high-dimensional gene regulatory network that exhibits oscillatory behavior, driven by a positive-negative feedback loop mimicking a circadian clock [Vilar et al., 2002].
 The model encompasses 9 species undergoing 18 reactions, parameterized by 15 reaction rate constants.
 Although the full system involves multiple interacting components, we focus on the inference of only three species (C, A, R).
-Detailed description is available in Appendix A.
+Detailed description is available in Appendix B.
 Unlike the SBI benchmarking problems, the true posterior distributions for these real-world models are unknown.
-To assess ConDiSim, we use well motivated parameter combinations from literature to generate corresponding synthetic observations (e.g., parameter combination resulting in stable oscillations in case of the genetic oscillator).
-These observations serve as the observed data for inference, enabling us to evaluate the inferred posterior by comparing its concentration around the true parameter value.
+To assess ConDiSim, we use well motivated parameter combinations from literature to generate corresponding synthetic observations (e.g., parameter combination where the genetic oscillator results in stable oscillations).
+Performing inference based on these observations enables us to evaluate the inferred posterior by comparing its concentration around the true parameter value.
 We also perform a posterior predictive check to validate the model.
@@
 Figure 5(b) illustrates both the true and generated trajectories for three selected species, obtained by randomly sampling three posterior sets and plotting the corresponding simulations.
-Initially, these trajectories align well with the true data, with slight deviations emerging over time due to the inherent stochastic nature of the model.
+The trajectories align well with the true data, with slight deviations emerging over time due to the inherent stochastic nature of the model, as expected.
 Figure 6 (Appendix) depicts multiple trajectories simulated using the true parameter combination - which closely match the inferred results.
 In both cases, the trajectories are very close to each other during initial time steps, and develop small deviations owing to model stochasticity as time steps evolve.
-The plots highlight the system’s inherent variability, and also validating the inferred solution.
+The plots highlight the system’s inherent variability, and also validate the inferred solution.
+From Figs.5(a), 9, we also see that ConDiSim is able to infer a more accurate, tighter posterior as compared to Simformer, highlighting its suitability for challenging inference problems involving stochastic, discrete-time observations.
 5 Discussion
-We analyze the impact of the hyperparameters — variance schedule ($\beta_{t}$) and diffusion steps ($T$), on ConDiSim’s performance in Section 5.1, with further discussion on overall scalability, robustness, and reliability in Sections 5.2, 5.4, and 5.5 respectively.
-This study provides practical insights for optimizing accuracy, efficiency, and reliability across various simulation-based inference tasks.
-5.1 Variance Schedule & Diffusion Steps
-We conducted an ablation study on the Two Moons test problem to examine the effects of diffusion steps ($T$) and variance schedules ($\beta_{t}$) on posterior approximation accuracy.
-Varying $T$ from 100 to 1000, we calculate the mean Classifier Two-Sample Test (C2ST) accuracy across five runs with a fixed simulation budget of 10,000 samples, while holding architecture and other hyperparameters constant.
-The study compared linear, cosine, and quadratic schedules to assess their impact under varying diffusion granularity, with $T$ controlling the refinement steps and $\beta_{t}$ setting the noise rate in the forward diffusion process.
-Figure 6: Mean C2ST score across diffusion steps for linear, cosine, and quadratic schedules on the Two Moons benchmark.
-Table 1: Comparison of C2ST scores across SBI benchmarks: (mean $\pm$ std.) at varying simulation budgets.
-Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 1.0 $\pm$0.0000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0000 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.$9$$\pm$0.0001 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.$9$$\pm$0.0003 0.$9$$\pm$0.0012 1.$0$$\pm$0.0002 FMPE 0.9713 $\pm$ 0.0117 0.9779 $\pm$ 0.0095 0.9387 $\pm$ 0.0377 0.9742 $\pm$ 0.0113 0.6661 $\pm$ 0.0307 0.6653 $\pm$ 0.0377 0.$9$$\pm$ 0.0178 0.$9$$\pm$ 0.0345 0.$9$$\pm$ 0.0018 ConDiSim 0.9722 $\pm$0.0368 0.9366 $\pm$0.02395 0.9347 $\pm$0.0368 0.6430 $\pm$0.0237 0.6155 $\pm$0.0317 0.6166 $\pm$0.0480 0.9996 $\pm$0.0002 0.9995 $\pm$0.0002 0.9996 $\pm$0.0002
-Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7474 $\pm$0.0364 0.6977 $\pm$0.0260 0.7234 $\pm$0.0599 0.9930 $\pm$0.0038 0.9916 $\pm$0.0046 0.9929 $\pm$0.0059 0.9989 $\pm$0.0004 0.9976 $\pm$0.0016 0.9988 $\pm$0.0010 0.9687 $\pm$0.0286 0.9436 $\pm$0.0203 0.9723 $\pm$0.0140 NPE 0.6118 $\pm$0.0155 0.5830 $\pm$0.0240 0.5818 $\pm$0.0158 0.5472 $\pm$0.0076 0.5238 $\pm$0.0112 0.5266 $\pm$0.0037 0.5562 $\pm$0.0147 0.5292 $\pm$0.0068 0.5256 $\pm$0.0066 0.8996 $\pm$0.0206 0.8502 $\pm$0.0174 0.8546 $\pm$0.0213 APT 0.5457 $\pm$0.0058 0.5425 $\pm$0.0124 0.5467 $\pm$0.0050 0.5367 $\pm$0.0038 0.5293 $\pm$0.0052 0.5258 $\pm$0.0047 0.5401 $\pm$0.0115 0.5302 $\pm$0.0069 0.5264 $\pm$0.0075 0.8164 $\pm$0.0295 0.6882 $\pm$0.0342 0.6473 $\pm$0.0343 Simformer 0.5166 $\pm$0.0210 0.5100 $\pm$0.0090 0.5126 $\pm$0.0070 0.6130 $\pm$0.0281 0.6226 $\pm$0.0398 0.5866 $\pm$0.0430 0.6501 $\pm$0.0519 0.6207 $\pm$0.0519 0.5780 $\pm$0.0211 0.7661 $\pm$0.0685 0.6201 $\pm$0.0490 0.7192 $\pm$0.0721 FMPE 0.7894 $\pm$ 0.0254 0.7872 $\pm$ 0.0049 0.5821 $\pm$ 0.0292 0.9605$\pm$ 0.0019 0.9716 $\pm$ 0.0023 0.9708 $\pm$ 0.0025 0.9218 $\pm$ 0.0122 0.9100 $\pm$ 0.0174 0.9082 $\pm$ 0.0178 0.9698 $\pm$ 0.0132 0.9686 $\pm$ 0.0139 0.9686 $\pm$ 0.0139 ConDiSim 0.6602 $\pm$0.0098 0.6181 $\pm$0.0237 0.6233 $\pm$0.0276 0.5809 $\pm$0.0210 0.5656 $\pm$0.0197 0.5505 $\pm$0.0056 0.6572 $\pm$0.0215 0.6468 $\pm$0.0364 0.6448 $\pm$0.0216 0.8691 $\pm$0.0132 0.8324 $\pm$0.0281 0.8210 $\pm$0.0188
-Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.7770 $\pm$0.0630 0.6634 $\pm$0.0503 0.7467 $\pm$0.1036 0.$9$$\pm$0.0001 0.$9$$\pm$0.0000 0.9997 $\pm$0.0005 0.9990 $\pm$0.0015 0.9912 $\pm$0.0161 0.9838 $\pm$0.0157 NPE 0.5749 $\pm$0.0331 0.5931 $\pm$0.0329 0.5411 $\pm$0.0195 0.9121 $\pm$0.0338 0.8988 $\pm$0.0306 0.8898 $\pm$0.0565 0.8378 $\pm$0.0396 0.9122 $\pm$0.0207 0.9405 $\pm$0.0125 APT 0.5189 $\pm$0.0080 0.5202 $\pm$0.0114 0.5147 $\pm$0.0125 0.8404 $\pm$0.0107 0.8328 $\pm$0.0156 0.8353 $\pm$0.0164 0.9433 $\pm$0.0076 0.9434 $\pm$0.0061 0.9457 $\pm$0.0068 Simformer 0.5273 $\pm$0.0229 0.5190 $\pm$0.0174 0.5098 $\pm$0.0084 0.6584 $\pm$0.0381 0.6131 $\pm$0.0225 0.6154 $\pm$0.0265 0.9378 $\pm$0.0053 0.9520 $\pm$0.0064 0.9478 $\pm$0.0049 FMPE 0.8051 $\pm$ 0.0506 0.7745 $\pm$ 0.0261 0.5706 $\pm$ 0.0209 0.9266 $\pm$ 0.0288 0.9118 $\pm$ 0.0376 0.8431 $\pm$ 0.0393 0.9573 $\pm$ 0.0013 0.8454 $\pm$ 0.0014 0.8101 $\pm$ 0.0020 ConDiSim 0.5291 $\pm$0.0143 0.5395 $\pm$0.0248 0.5242 $\pm$0.0063 0.6589 $\pm$0.0421 0.5892 $\pm$0.0100 0.6146 $\pm$0.0371 0.8324 $\pm$0.0332 0.8083 $\pm$0.0157 0.8043 $\pm$0.0138
-Table 2: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.).
-Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 117.67 $\pm$2.43 164.90 $\pm$57.10 249.50 $\pm$105.05 118.62 $\pm$0.18 179.69 $\pm$49.96 253.48 $\pm$113.38 227.16 $\pm$1.44 241.24 $\pm$7.58 250.47 $\pm$2.22 NPE 7.04 $\pm$0.21 14.21 $\pm$0.04 20.95 $\pm$0.37 12.17 $\pm$0.17 24.96 $\pm$0.64 37.43 $\pm$0.22 20.10 $\pm$5.45 29.62 $\pm$3.74 44.92 $\pm$9.24 APT 90.31 $\pm$8.99 186.17 $\pm$13.79 268.28 $\pm$32.88 101.39 $\pm$7.89 197.30 $\pm$9.32 267.38 $\pm$21.93 40.15 $\pm$3.90 73.74 $\pm$6.25 112.51 $\pm$2.51 Simformer 3.96 $\pm$0.04 7.36 $\pm$0.11 10.73 $\pm$0.25 17.87 $\pm$0.18 35.05 $\pm$0.52 52.81 $\pm$0.91 4.24 $\pm$0.29 7.67 $\pm$0.38 10.68 $\pm$0.45 FMPE 2.63 $\pm$ 0.05 4.73$\pm$ 1.11 10.655 $\pm$ 0.04 2.65 $\pm$ 0.05 5.75 $\pm$ 1.43 7.64 $\pm$ 1.38 2.78 $\pm$ 0.15 6.39 $\pm$ 0.19 8.94 $\pm$ 2.94 CondiSim 0.84 $\pm$0.28 1.682 $\pm$0.09 2.835 $\pm$0.70 0.31 $\pm$0.11 0.60 $\pm$0.14 0.88 $\pm$0.17 0.87 $\pm$0.16 0.94 $\pm$0.13 1.59 $\pm$0.24
-Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 FMPE 2.76 $\pm$ 0.08 3.08 $\pm$ 1.76 12.97 $\pm$ 0.87 2.44 $\pm$ 0.44 4.42 $\pm$ 1.37 8.05 $\pm$ 1.38 3.63 $\pm$ 0.42 6.97 $\pm$ 0.03 9.55 $\pm$ 0.72 CondiSim 1.31 $\pm$0.07 1.83 $\pm$0.32 2.08 $\pm$0.34 0.43 $\pm$0.04 0.62 $\pm$0.08 0.93 $\pm$0.17 0.60 $\pm$0.11 1.15 $\pm$0.04 1.41 $\pm$0.08
-Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 114.07 $\pm$2.11 111.30 $\pm$0.28 137.46 $\pm$19.16 48.19 $\pm$0.24 48.70 $\pm$0.84 49.12 $\pm$0.49 43.27 $\pm$0.36 66.43 $\pm$5.70 69.10 $\pm$0.15 162.72 $\pm$14.73 196.19 $\pm$119.42 99.39 $\pm$5.63 NPE 19.56 $\pm$2.47 42.88 $\pm$4.54 64.03 $\pm$6.51 12.03 $\pm$0.80 24.69 $\pm$1.63 38.79 $\pm$1.23 16.55 $\pm$0.85 30.90 $\pm$2.77 49.42 $\pm$1.49 42.97 $\pm$6.35 97.70 $\pm$6.66 117.61 $\pm$4.12 APT 76.27 $\pm$7.02 167.88 $\pm$15.99 240.13 $\pm$22.74 70.47 $\pm$0.88 135.38 $\pm$5.75 203.27 $\pm$5.97 78.61 $\pm$2.11 158.57 $\pm$9.32 219.44 $\pm$11.83 93.40 $\pm$1.26 150.29 $\pm$11.10 216.30 $\pm$21.38 Simformer 3.84 $\pm$0.04 7.15 $\pm$0.09 10.48 $\pm$0.21 18.18 $\pm$0.13 35.46 $\pm$0.17 55.46 $\pm$1.13 17.89 $\pm$0.23 33.47 $\pm$0.40 51.24 $\pm$0.89 4.23 $\pm$0.01 7.67 $\pm$0.13 11.32 $\pm$0.40 FMPE 2.16$\pm$ 0.08 5.36 $\pm$ 0.78 8.31 $\pm$ 1.32 1.78 $\pm$ 0.05 4.40 $\pm$ 0.78 5.94 $\pm$ 4.31 2.15 $\pm$ 0.22 4.30 $\pm$ 3.86 8.71 $\pm$ 2.12 1.36 $\pm$ 0.14 3.71 $\pm$ 2.57 6.34 $\pm$ 2.33 ConDiSim 0.74 $\pm$0.18 1.77 $\pm$0.41 2.31 $\pm$0.59 0.76 $\pm$0.09 1.95 $\pm$0.59 3.46 $\pm$0.70 0.91 $\pm$0.20 2.34 $\pm$0.37 3.02 $\pm$0.69 1.13 $\pm$0.08 2.61 $\pm$0.80 2.82 $\pm$0.36
-As shown in Figure 6, lower $T$ values generally yield better C2ST accuracy, potentially due to reduced cumulative noise, which preserves crucial posterior structure.
-The quadratic schedule outperforms others, possibly because its gradual noise addition captures complex structures more effectively without early-stage noise overload.
-Increasing $T$ leads to a plateau and slight decline in accuracy, suggesting diminishing returns and rising instability, with all schedules converging in performance at high $T$ values.
+In this section, we compare ConDiSim against established SBI methods (Section 5.1) and analyze its behavior along three practical axes: scalability with problem dimensionality (Section 5.2), robustness to simulator noise or misspecification (Section 5.3), and reliability in capturing posterior uncertainty (Section 5.4).
+This analysis provides practical guidance for deploying diffusion-based inference methods by identifying regimes where ConDiSim excels and scenarios where alternative approaches may be more suitable.
+5.1 ConDiSim vs. Other SBI Methods
+Tables 1 and 4 show that diffusion-based methods (Simformer, ConDiSim) often achieve higher posterior accuracy than flow-based (NPE, APT, FMPE) and GAN-based (GATSBI) approaches, particularly on problems with multimodal or highly nonlinear posteriors.
+Simformer attains the best overall performance on several benchmarks (Table 4), owing to its transformer architecture with cross-attention conditioning; however, this comes at the cost of substantially longer training times (Table 2).
+In contrast, ConDiSim uses a lightweight FiLM-based conditioning mechanism that avoids costly attention computations, achieving competitive accuracy with significantly lower training overhead - while its iterative denoising incurs no practical slowdown in inference.
+NPE and APT remain strong baselines on simpler or lower-dimensional tasks, where their one-shot or sequential flow-based strategies are sufficient.
+GATSBI, meanwhile, is hampered by the well-known instabilities of adversarial training.
+We also note from Table 1 that ConDiSim outperforms other methods on problems involving noisy, discrete observations, such as Bernoulli GLM, Bernoulli GLM Raw and SIR, leveraging FiLM.
+This makes the proposed approach well-suited for SBI problems in computational biology where parameter inference of stochastic models with time-series observations is often encountered (e.g., the genetic oscillator test problem, as noted before).
+Table 1: C2ST scores across SBI benchmarks (mean $\pm$ std.) at varying simulation budgets; best in bold, second-best underlined.
+Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 0.9714 $\pm$0.0144 0.9468 $\pm$0.0157 0.9359 $\pm$0.0127 0.9354 $\pm$0.0093 0.9114 $\pm$0.0082 0.9065 $\pm$0.0105 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0000 APT 0.9823 $\pm$0.0055 0.9451 $\pm$0.0158 0.9130 $\pm$0.0301 0.8455 $\pm$0.0116 0.8596 $\pm$0.0152 0.8461 $\pm$0.0135 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 GATSBI 0.9995 $\pm$0.0007 0.9978 $\pm$0.0016 0.9994 $\pm$0.0004 0.9998 $\pm$0.0001 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 0.9999 $\pm$0.0001 0.9999 $\pm$0.0000 1.0 $\pm$0.0000 Simformer 0.9198 $\pm$0.0451 0.8814 $\pm$0.0433 0.8717 $\pm$0.0453 0.9814 $\pm$0.0106 0.8600 $\pm$0.0452 0.7750 $\pm$0.0473 0.9999 $\pm$0.0003 0.9999 $\pm$0.0012 1.0000 $\pm$0.0002 FMPE 0.9713 $\pm$ 0.0117 0.9779 $\pm$ 0.0095 0.9387 $\pm$ 0.0377 0.9742 $\pm$ 0.0113 0.6661 $\pm$ 0.0307 0.6653 $\pm$ 0.0377 0.9999 $\pm$ 0.0178 0.9999 $\pm$ 0.0345 0.9999 $\pm$ 0.0018 ConDiSim 0.9456 $\pm$0.0245 0.9222 $\pm$0.02609 0.8832 $\pm$0.08519 0.6673 $\pm$0.0668 0.6099 $\pm$0.0272 0.6041 $\pm$0.0237 0.9995 $\pm$0.0001 0.9995 $\pm$0.0005 0.9995 $\pm$0.0003
+[…]
+[…]
+Table 2: Training time (in minutes) for different methods across simulation budgets (mean $\pm$ std. dev.); best in bold, second-best underlined.
+[…]
+Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 21.92 $\pm$2.41 43.92 $\pm$2.58 64.38 $\pm$7.13 25.39 $\pm$2.09 61.43 $\pm$2.61 85.94 $\pm$6.55 80.67 $\pm$14.61 160.92 $\pm$42.82 188.04 $\pm$29.46 APT 26.30 $\pm$0.36 49.96 $\pm$2.73 74.31 $\pm$2.54 21.87 $\pm$1.28 40.00 $\pm$1.37 60.94 $\pm$3.00 45.52 $\pm$4.85 96.15 $\pm$13.77 144.33 $\pm$18.33 GATSBI 74.37 $\pm$2.59 100.59 $\pm$2.39 102.30 $\pm$0.51 114.76 $\pm$1.19 115.22 $\pm$1.43 144.10 $\pm$24.15 78.91 $\pm$6.32 75.73 $\pm$0.70 74.99 $\pm$0.20 Simformer 15.98 $\pm$1.22 31.98 $\pm$2.34 47.85 $\pm$1.86 16.96 $\pm$0.51 32.46 $\pm$1.89 48.01 $\pm$1.44 5.0 $\pm$0.34 9.23 $\pm$0.54 13.64 $\pm$0.21 FMPE 2.76 $\pm$ 0.08 3.08 $\pm$ 1.76 12.97 $\pm$ 0.87 2.44 $\pm$ 0.44 4.42 $\pm$ 1.37 8.05 $\pm$ 1.38 3.63 $\pm$ 0.42 6.97 $\pm$ 0.03 9.55 $\pm$ 0.72 CondiSim 2.02 $\pm$0.27 5.16 $\pm$0.84 8.44 $\pm$1.02 3.22 $\pm$0.72 7.22 $\pm$0.50 10.67 $\pm$0.69 2.42 $\pm$0.21 7.19 $\pm$0.97 11.19 $\pm$2.31
+[…]
 5.2 Scalability
-ConDiSim’s scalability was extensively evaluated across diverse benchmark problems, spanning from low-dimensional (Two Moons, GMM, SIR; 2 parameters) to higher-dimensional scenarios (Gaussian Linear, Bernoulli GLM; 10 parameters, HH Model; 7 parameters, genetic oscillator; 15 parameters).
-From the C2ST metric values in Table 1, and the posterior plots spanning from Figure 2(a) to Figure 13, it can be seen that ConDiSim maintained the posterior approximation accuracy with increasing problem dimensionality and complexity.
-Model training times in Table 2 also show that ConDiSim is computationally efficient to train and scales favorably with problem complexity, making it suitable for high-dimensional SBI tasks.
-Its ability to handle larger parameter spaces without significant loss of performance is a key advantage, as evidenced in the genetic oscillator test problem involving 15 parameters.
-5.3 Comparative Evaluation of different Inference Methods
-Tables 1 and 4 indicate that diffusion models (Simformer, ConDiSim) generally outperform flow-based (NPE, APT) and GAN-based (GATSBI) approaches in accuracy, owing to their ability to model complex posteriors.
-Simformer achieves more accurate posterior approximations in most test problems (also see Table 4) owing to its complex architecture, however it also entails longer training times due to the involved complexity (Table 2).
-In contrast, ConDiSim’s simplified conditioning mechanism and architecture reduces training complexity, and its iterative denoising process does not impede inference speed in our experiments.
-The known challenges of adversarial training affect GATSBI, while APT’s sequential nature increases its computational cost relative to NPE.
+We evaluate ConDiSim across a spectrum of SBI benchmarks, ranging from low-dimensional problems (Two Moons, GMM, SIR; 2 parameters) to higher-dimensional settings (Gaussian Linear, Bernoulli GLM; 10 parameters, HH Model; 7 parameters, genetic oscillator; 15 parameters).
+As shown by C2ST scores in Table 1 and posterior visualizations (Figures 2(a)13), ConDiSim maintains high posterior fidelity even as dimensionality and model complexity increase.
+Training times in Table 2 further confirm its computational efficiency, with favorable scaling relative to problem size.
+Notably, the 15-dimensional genetic oscillator highlights ConDiSim’s architectural efficiency and effectiveness, as it preserves inference accuracy without suffering substantial degradation owing to a relatively large parameter space and the noisy, stochastic nature of the simulator.
 5.3 Robustness
-Robustness to observational noise is a critical factor in the practical application of simulation-based inference (SBI) methods.
-To rigorously evaluate this aspect, we conducted experiments on two challenging tasks: the SLCP Distractors and the Bernoulli GLM Raw.
-The SLCP Distractors task introduces 92 non-informative variables to the observations, significantly increasing the inference complexity.
-In contrast, the Bernoulli GLM Raw task utilizes raw data instead of well-behaved summary statistics, presenting a more realistic scenario often encountered in real-world applications.
-The results for the SLCP Distractors and Bernoulli GLM RAW, illustrated in Table 1 and Figures 3(b),12(b) demonstrate ConDiSim’s robustness to both observational noise.
-Such resilience is important as it indicates ConDiSim’s capacity to effectively handle real-world data, which is often characterized by measurement errors and inherent uncertainties.
+In many real-world problems, observational data are inherently imperfect and often contaminated by measurement noise, confounded by irrelevant features, or available only in raw, unprocessed form.
+Consequently, it is essential to evaluate how inference methods perform under such adverse conditions.
+We assess ConDiSim’s robustness on two challenging benchmarks: SLCP Distractors, where 92 uninformative variables are added to the observation vector to simulate high-dimensional nuisance features, and Bernoulli GLM Raw, which operates directly on raw binary observations rather than relying on handcrafted summary statistics, emulating realistic scenarios with noisy, unprocessed measurements.
+The genetic oscillator test problem is also noisy, simulated using the Stochastic Simulation Algorithm (SSA) [Gillespie, 2007].
+As shown in Table 1 and Figures 3(b)12(b), ConDiSim maintains accurate posterior recovery across both settings, demonstrating resilience to observation noise and feature redundancy that demonstrates its suitability for practical scientific applications where data are high-dimensional, unprocessed, or corrupted by measurement uncertainty.
 5.4 Reliability
-To further validate the learned inference model, we employed Simulation-Based Calibration (SBC) as a rigorous diagnostic for assessing the quality of approximate posterior distributions.
-The empirical cumulative distribution function (ECDF) plots demonstrate ConDiSim’s ability to produce accurate and well-calibrated posteriors, with rank statistics closely aligning along the diagonal—indicative of unbiased and reliable inference.
-This alignment supports the statistical validity and robustness of ConDiSim’s inferences across diverse and challenging scenarios.
+Beyond predictive accuracy, it is crucial to verify that the learned posteriors are statistically well-calibrated, meaning that the reported uncertainty faithfully captures the true variability of the inference problem.
+To assess calibration, we use Simulation-Based Calibration (SBC) [Talts et al., 2018], which is a standard diagnostic that evaluates whether the rank statistics of posterior samples follow their expected uniform distribution.
+The empirical cumulative distribution function (ECDF) differences shown in Figures 2(b)15 remain within the expected bounds, with fluctuations centered around zero, indicating that ConDiSim produces unbiased and well-calibrated posteriors across benchmarks.
+Such calibration is essential for trustworthy inference in scientific applications, where both accurate point estimates and reliable uncertainty quantification are required.
 6 Conclusion
-In this work, we introduced ConDiSim, a denoising conditional diffusion model for simulation-based inference.
-We demonstrated ConDiSim’s ability to effectively capture complex posterior distributions efficiently on 10 benchmark problems and 2 real-world test problems.
-ConDiSim performs comparably to the state-of-the-art while being substantially faster to train.
-As part of an ablation study, we also explored the significance of the variance schedule and the number of diffusion steps in influencing model performance.
-ConDiSim offers a flexible and efficient approach to simulation-based inference, with potential applications across various scientific and engineering domains.
-Future work includes exploring adaptive variance schedules, integrating domain-specific knowledge into the model, and extending ConDiSim towards learning both the likelihood and posterior distributions.
+We introduce ConDiSim, a DDPM-based conditional diffusion model for simulation-based inference.
+Across ten benchmark tasks and two real-world problems, ConDiSim effectively handles ill-posed inference, multimodal and high-dimensional posteriors, and demonstrates robustness even with raw, unprocessed observations, or in the presence of noise and distractors.
+It also adapts naturally to diverse data modalities, while offering substantially faster training time than competing methods.
+We demonstrate that ConDiSim is particularly suited for inference of stochastic simulators with time series output, where it outperforms other methods and delivers accuracy, well-calibrated posterior estimates, and fast amortized inference.
 Acknowledgments
@@
 2023-05167 and 2023-05593 respectively.
-Appendix A SBI Problem Description
+Appendix A Connection between ConDiSim and Score based Diffusion Model
+In this section, we establish a formal theoretical connection between ConDiSim, a DDPM based diffusion model, and the score based diffusion models.
+We demonstrate its equivalence to a variance preserving stochastic differential equation (SDE) by deriving the continuous time limit of ConDiSim’s discrete forward process, providing a unified perspective on diffusion based generative modeling.
+A.1 Forward Process and Continuous-Time Limit
+ConDiSim employs a discrete time forward process, defined as:
+$\bm{\theta}_{t}=\sqrt{1-\beta_{t}}\,\bm{\theta}_{t-1}+\sqrt{\beta_{t}}\,\bm{\epsilon},\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}),$ (17)
+where $\bm{\theta}_{t}$ represents the state at time step $t$, $\beta_{t}$ is the noise schedule, and $\bm{\epsilon}$ is standard Gaussian noise.
+To convert this discrete time process to continuous time, we consider the limit as the number of time steps $T\to\infty$ and define a small time step $\Delta t=\frac{1}{T}\to 0$.
+We introduce a continuous noise schedule $\beta(t)$ such that $\beta_{t}\approx\beta(t)\Delta t$.
+Using a first-order Taylor expansion for small $\beta_{t}$, we approximate $\sqrt{1-\beta_{t}}\approx 1-\frac{1}{2}\beta(t)\Delta t$.
+Substituting into the forward process:
+$\bm{\theta}_{t}=\left(1-\frac{1}{2}\beta(t)\Delta t\right)\bm{\theta}_{t-1}+\sqrt{\beta(t)\Delta t}\,\bm{\epsilon},$ (18)
+Defining the increment $\Delta\bm{\theta}_{t}=\bm{\theta}_{t}-\bm{\theta}_{t-1}$, and noting that the noise term $\sqrt{\beta(t)\Delta t}\,\bm{\epsilon}$ corresponds to a Wiener process increment $\Delta\mathbf{W}\sim\mathcal{N}(\mathbf{0},\Delta t\mathbf{I})$, we rewrite the equation as:
+$\Delta\bm{\theta}_{t}=-\frac{1}{2}\beta(t)\bm{\theta}_{t-1}\Delta t+\sqrt{\beta(t)}\,\Delta\mathbf{W}.$ (19)
+In the continuous-time limit ($\Delta t\to 0$), we replace $\Delta\bm{\theta}_{t}\to d\bm{\theta}_{t}$, $\Delta t\to dt$, $\Delta\mathbf{W}\to d\mathbf{w}_{t}$, and $\bm{\theta}_{t-1}\simeq\bm{\theta}_{t}$, yielding the following stochastic differential equation (SDE):
+$d\bm{\theta}_{t}=-\frac{1}{2}\beta(t)\bm{\theta}_{t}\,dt+\sqrt{\beta(t)}\,d\mathbf{w}_{t},$ (20)
+where $\mathbf{w}_{t}$ denotes standard Brownian motion and $d\mathbf{w}_{t}\sim\mathcal{N}(0,dt\,\mathbf{I})$.
+Comparing Equation 20 with the general form of the stochastic differential equations (SDEs):
+$d\bm{\theta}_{t}=\mathbf{f}(\bm{\theta}_{t},t)\,dt+g(t)\,d\mathbf{w}_{t},$ (21)
+where, $\mathbf{f}(\bm{\theta}_{t},t)$ is the deterministic drift term and $g(t)$ is the scalar valued diffusion coefficient.
+Equation (20) is a special case of the general SDE, with $\mathbf{f}(\bm{\theta}_{t},t)=-\frac{1}{2}\beta(t)\bm{\theta}_{t}$ and $g(t)=\sqrt{\beta(t)}$; this is known as the Variance-Preserving SDE (VP-SDE) [Song et al., 2021] and corresponds to an Ornstein–Uhlenbeck (OU) process.
+The solution to the VP-SDE can be derived using Itô calculus and is given by:
+$\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon},\quad\text{where }\alpha_{t}=\exp\left(-\int_{0}^{t}\beta(s)\,ds\right),\quad\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$ (22)
+This expression represents the marginal distribution $q(\bm{\theta}_{t}\mid\bm{\theta}_{0})$ under the continuous diffusion process and serves as the continuous time analog of the discrete DDPM forward process:
+$\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon},\quad\alpha_{t}=\prod_{i=1}^{t}(1-\beta_{i}).$ (23)
+Thus, the VP-SDE generalizes the discrete Markov chain of ConDiSim to continuous time.
+A.2 Reverse Diffusion Process and Score Matching
+Given a general forward SDE as in Equation 21, the associated reverse-time SDE describes the dynamics when running the process backward in time (from $t=T$ to $t=0$) and corresponds to the reverse diffusion process of ConDiSim.
+Following [Anderson, 1982], the general form of the reverse SDE is:
+$d\bm{\theta}_{t}=\left[\mathbf{f}(\bm{\theta}_{t},t)-g^{2}(t)\,\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right]dt+g(t)\,d\mathbf{w}_{t},$ (24)
+where $q_{t}(\bm{\theta}_{t})$ is the marginal probability density of $\bm{\theta}_{t}$ at time $t$ and $d\mathbf{w}_{t}$ is Brownian motion in reverse time.
+For the VP-SDE (Equation (20)), the reverse SDE becomes:
+$d\bm{\theta}_{t}=\left[-\frac{1}{2}\beta(t)\,\bm{\theta}_{t}-\beta(t)\,\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right]dt+\sqrt{\beta(t)}\,d\mathbf{w}_{t}.$ (25)
+In this formulation, all terms are known except for the score function $s(\bm{\theta}_{t},t)=\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})$, which is crucial for simulating the reverse process.
+Approximating this score is a standard optimization problem known as score matching, which is well-studied in probabilistic sampling, especially in the context of Langevin dynamics.
+However, since the true score depends on the intractable marginal $q_{t}(\bm{\theta}_{t})$, we instead introduce a parameterized approximation $s_{\phi}(\bm{\theta}_{t},t)$ and learn it by minimizing the Fisher divergence (explicit score matching loss):
+$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{t}(\bm{\theta}_{t})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)-\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t})\right\|^{2}\right].$ (26)
+To address the intractability of the true score, several methods have been developed, including implicit score matching [Hyvärinen and Dayan, 2005], sliced score matching[Song et al., 2020], and denoising score matching [Vincent, 2011].
+To align the training objective with ConDiSim, we adopt denoising score matching (DSM), which leverages the tractable conditional score $\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}\mid\bm{\theta}_{0})$.
+The corresponding objective is given by:
+$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{0}(\bm{\theta}_{0})}\;\mathbb{E}_{q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)-\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})\right\|^{2}\right].$ (27)
+Given the closed-form solution for the forward process in Equation (22), the score function $\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}\mid\bm{\theta}_{0})$ becomes tractable and can be written as:
+$\nabla_{\bm{\theta}_{t}}\log q_{t}(\bm{\theta}_{t}|\bm{\theta}_{0})=-\frac{\bm{\theta}_{t}-\sqrt{\alpha(t)}\,\bm{\theta}_{0}}{1-\alpha(t)}.$ (28)
+Reparameterizing $\bm{\theta}_{t}=\sqrt{\alpha_{t}}\,\bm{\theta}_{0}+\sqrt{1-\alpha_{t}}\,\bm{\epsilon}$ with $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$, the DSM objective becomes:
+$\phi^{*}=\arg\min_{\phi}\;\mathbb{E}_{t}\;\mathbb{E}_{q_{0}(\bm{\theta}_{0})}\;\mathbb{E}_{\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})}\left[\left\|s_{\phi}(\bm{\theta}_{t},t)+\frac{1}{\sqrt{1-\alpha_{t}}}\,\bm{\epsilon}\right\|^{2}\right].$ (29)
+This objective aligns closely with the ConDiSim training objective, where the score is approximated as:
+$s_{\phi}(\bm{\theta}_{t},t)\approx-\frac{1}{\sqrt{1-\alpha_{t}}}\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t},t),$ (30)
+with $\hat{\bm{\epsilon}}_{\phi}(\bm{\theta}_{t},t)$ being the predicted noise.
+This equivalence demonstrates that ConDiSim’s training objective is a rescaled form of the DSM objective, unifying the discrete DDPM framework with continuous-time score-based diffusion models.
+This analysis reveals that, although ConDiSim is formulated as a discrete DDPM, it is fundamentally equivalent to a score-based diffusion model in the continuous-time limit.
+The VP-SDE provides a principled framework for analyzing ConDiSim’s dynamics: the forward diffusion process can be understood through the Fokker–Planck equation, which characterizes convergence properties of the marginal distributions, while the reverse diffusion process connects naturally to annealed Langevin dynamics.
+Appendix B SBI Problem Description
 We present the general setup for simulation-based inference (SBI), which is employed across various benchmarking problems to estimate the posterior distribution $p(\bm{\theta}|\mathbf{y})$ over parameters $\bm{\theta}$, given observed data $\mathbf{y}$.
@@
 High-dimensional spaces often present greater challenges in accurately estimating the posterior distribution.
-A.1 Two Moons
+B.1 Two Moons
 This parameter inference task is designed to test the performance of inference methods under multi-modal conditions, featuring a posterior distribution with both global (bi-modal) and local (crescent-shaped) characteristics.
-  • Prior: $\bm{\theta}\sim\mathcal{U}(-1,1)$
+  • Prior: $\bm{\theta}\sim\mathcal{U}(-1,1)$.
   • Simulator: $\mathbf{y}|\bm{\theta}=r\begin{bmatrix}\cos(\alpha)\\ \sin(\alpha)\end{bmatrix}+0.25\begin{bmatrix}-|\theta_{1}+\theta_{2}|/\sqrt{2}\\ (-\theta_{1}+\theta_{2})/\sqrt{2}\end{bmatrix}$, where $\alpha\sim\mathcal{U}(-\pi/2,\pi/2)$ and $r\sim\mathcal{N}(0.1,0.01^{2})$.
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
-A.2 Gaussian Mixture
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$.
+B.2 Gaussian Mixture
 This task challenges inference methods to estimate the shared mean of a mixture of Gaussian distributions with distinct covariances, thereby assessing their performance in handling multi-modal and overlapping data distributions.
-  • Prior: $\bm{\theta}\sim\mathcal{U}(-10,10)$
-  • Simulator: $\mathbf{y}|\bm{\theta}=0.5\left(\mathcal{N}(\bm{\theta},\mathbf{I})+\mathcal{N}(\bm{\theta},0.01\mathbf{I})\right)$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$
-A.3 Gaussian Linear
+  • Prior: $\bm{\theta}\sim\mathcal{U}(-10,10)$.
+  • Simulator: $\mathbf{y}|\bm{\theta}=0.5\left(\mathcal{N}(\bm{\theta},\mathbf{I})+\mathcal{N}(\bm{\theta},0.01\mathbf{I})\right)$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{2}$.
+B.3 Gaussian Linear
 This task involves estimating the mean vector $\bm{\theta}$ of a multivariate Gaussian distribution with a fixed and known covariance matrix.
@@
 This conjugacy makes the task an ideal benchmark for evaluating inference methods, as it allows for a clear assessment of their ability to accurately capture known posterior distributions.
-  • Prior: $\bm{\theta}\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$
-  • Simulator: Generates data $\mathbf{y}$ using $\mathbf{y}|\bm{\theta}\sim\mathcal{N}(\bm{\theta},0.1\mathbf{I})$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
-A.4 Gaussian Linear Uniform
+  • Prior: $\bm{\theta}\sim\mathcal{N}(\mathbf{0},0.1\mathbf{I})$.
+  • Simulator: Generates data $\mathbf{y}$ using $\mathbf{y}|\bm{\theta}\sim\mathcal{N}(\bm{\theta},0.1\mathbf{I})$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$.
+B.4 Gaussian Linear Uniform
 This task is a variant of the Gaussian Linear problem, where the mean parameter $\bm{\theta}$ is assigned a uniform prior distribution $\bm{\theta}\sim\mathcal{U}(-1,1)$ instead of a Gaussian prior.
@@
 This modification tests the inference method’s ability to handle non-Gaussian priors and to accurately capture the resulting posterior distributions.
-A.5 SLCP
+B.5 SLCP
 The SLCP task presents a more challenging inference scenario where, despite the simplicity of the likelihood function, the posterior distribution exhibits complex, non-linear dependencies among parameters.
@@
 This setup assesses the inference method’s capability to accurately capture intricate posterior structures arising from non-linear relationships between parameters and observations.
-  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
+  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$.
   • Simulator: Outputs data $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{4})$, where:
-    $y_{i}\sim\mathcal{N}\left(\begin{bmatrix}\theta_{1}\\ \theta_{2}\end{bmatrix},\begin{bmatrix}\theta_{3}^{2}&\tanh(\theta_{5})\theta_{3}\theta_{4}\\ \tanh(\theta_{5})\theta_{3}\theta_{4}&\theta_{4}^{2}\end{bmatrix}\right)$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{8}$
-A.6 SLCP with Distractors
+[…]
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{8}$.
+B.6 SLCP with Distractors
 An extension of the SLCP task that incorporates additional non-informative dimensions (distractors) into the observed data.
 This setup evaluates the robustness of inference methods when faced with high-dimensional data containing redundant information.
-  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$
+  • Prior: $\bm{\theta}\sim\mathcal{U}(-3,3)$.
   • Simulator: Outputs data $y|\theta=(y_{1},\ldots,y_{100})$, where a permutation function $p(y)$ reorders $y$ with a fixed random permutation.
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{100}$
-A.7 Bernoulli GLM
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{5},\mathbf{y}\in\mathbb{R}^{100}$.
+B.7 Bernoulli GLM
 This task involves performing inference in a Generalized Linear Model (GLM) with Bernoulli-distributed observations and a 10-dimensional parameter vector $\bm{\theta}$.
@@
 This setup evaluates the inference method’s ability to handle discrete data and accurately capture the posterior distribution in models with latent smoothness constraints.
-  • Prior: $\beta\sim\mathcal{N}(\mathbf{0},2)$, $f\sim\mathcal{N}(\mathbf{0},(\mathbf{F}^{\top}\mathbf{F})^{-1})$
-  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$, with $\eta(z)=\frac{\exp(z)}{1+\exp(z)}$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$
-  • Fixed Parameters: $T=100$
-A.8 Bernoulli GLM Raw
+  • Prior: $\beta\sim\mathcal{N}(\mathbf{0},2)$, $f\sim\mathcal{N}(\mathbf{0},(\mathbf{F}^{\top}\mathbf{F})^{-1})$.
+  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$, with $\eta(z)=\frac{\exp(z)}{1+\exp(z)}$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{10}$.
+  • Fixed Parameters: $T=100$.
+B.8 Bernoulli GLM Raw
 This task is similar to the Bernoulli GLM described above but utilizes raw binary observations instead of sufficient statistics.
@@
 This scenario tests the method’s capability to accurately infer parameters in the presence of high-dimensional, discrete, and unprocessed data, challenging its robustness and flexibility.
-  • Prior: Same as Bernoulli GLM
-  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{100}$
-  • Fixed Parameters: $T=100$
-A.9 SIR Model
+  • Prior: Same as Bernoulli GLM.
+  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{100})$, where $y_{i}\sim\text{Bern}(\eta(\mathbf{V}_{i}^{\top}f+\beta))$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{10},\mathbf{y}\in\mathbb{R}^{100}$.
+  • Fixed Parameters: $T=100$.
+B.9 SIR Model
 The SIR model is an epidemiological framework that describes the number of individuals in three compartments: Susceptible ($S$), Infectious ($I$), and Recovered ($R$).
@@
 $\bm{\theta}=\begin{pmatrix}\beta\\ \gamma\end{pmatrix}\sim\text{LogNormal}\left(\begin{pmatrix}\log(0.4)\\ \log(1/8)\end{pmatrix},\begin{pmatrix}0.5\\ 0.2\end{pmatrix}\right)$.
-  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\mathcal{B}(1000,\frac{I}{N})$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{10}$
-  • Fixed Parameters: Population size $N=1,000,000$, task duration $T=160$
-A.10 Lotka-Volterra
+  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\mathcal{B}(1000,\frac{I}{N})$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{2},\mathbf{y}\in\mathbb{R}^{10}$.
+  • Fixed Parameters: Population size $N=1,000,000$, task duration $T=160$.
+B.10 Lotka-Volterra
 The Lotka-Volterra (LV) model is a classical ecological framework that describes the interaction between prey ($X$) and predator ($Y$) populations. The dynamics are expressed as:
@@
 $\bm{\theta}=\begin{pmatrix}\alpha\\ \beta\\ \gamma\\ \delta\end{pmatrix}\sim\text{LogNormal}\left(\begin{pmatrix}-0.125\\ -3\\ -0.125\\ -3\end{pmatrix},\begin{pmatrix}0.5\\ 0.5\\ 0.5\\ 0.5\end{pmatrix}\right)$.
-  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\text{LogNormal}(\log(X),0.1)$ and $y_{2i}\sim\text{LogNormal}(\log(Y),0.1)$
-  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{4},\mathbf{y}\in\mathbb{R}^{20}$
-  • Fixed Parameters: Task duration $T=20$, initial conditions $(X(0),Y(0))=(30,1)$
-A.11 Hodgkin-Huxley Model
+  • Simulator: $\mathbf{y}|\bm{\theta}=(y_{1},\ldots,y_{10})$, where $y_{i}\sim\text{LogNormal}(\log(X),0.1)$ and $y_{2i}\sim\text{LogNormal}(\log(Y),0.1)$.
+  • Dimensionality: $\bm{\theta}\in\mathbb{R}^{4},\mathbf{y}\in\mathbb{R}^{20}$.
+  • Fixed Parameters: Task duration $T=20$, initial conditions $(X(0),Y(0))=(30,1)$.
+B.11 Hodgkin-Huxley Model
 The Hodgkin-Huxley (HH) model describes the generation and propagation of action potentials in neurons by modeling the membrane voltage $V(t)$ as influenced by ionic currents through voltage-gated sodium ($\text{Na}^{+}$), potassium ($\text{K}^{+}$), and leak channels.
@@
   • Dimensionality: The parameter space is $\bm{\theta}\in\mathbb{R}^{7}$, and the output summary statistics $\mathbf{y}\in\mathbb{R}^{8}$.
-[…]
-A.12 High-Dimensional Genetic Oscillator
+  • Fixed Parameters: The initial membrane voltage is $V_{0}=-65.0\,\text{mV}$, with gating variables initialized to their steady-state values at $V_{0}$.
+Simulations use a time resolution of $5000\,\text{steps}$ over $200\,\text{ms}$ simulation window.
+B.12 Genetic Oscillator
 The genetic oscillator model describes a biochemical reaction network involving $9$ species governed by $18$ reactions, parameterized by $15$ reaction rate constants.
 This network is based on a positive–negative feedback loop, where the activator protein $A$ upregulates transcription of both its gene and the repressor protein $R$.
-The repressor $R$ sequesters $A$ by forming a complex $C$, thereby inhibiting further activation [Vilar et al., 2002].
+The repressor $R$ sequesters $A$ by forming a complex $C$, thereby inhibiting further activation [Vilar et al., 2002, Akesson et al., 2022].
 The species in the network are:
@@
   • $M_{A},M_{R}$: mRNA transcribed from activator ($M_{A}$) and repressor genes ($M_{R}$).
-  • $A$: Activator protein.
+  • AA: Activator protein.
   • $R$: Repressor protein.
   • $C$: Complex formed by $A$ and $R$.
-The network includes the following key reactions:
-$D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},$ $D_{R}^{*}$ $\xrightarrow{\theta_{R}}D_{R},$
-$D_{A},A$ $\xrightarrow{\gamma_{A}}D_{A}^{*},$ $D_{R},A$ $\xrightarrow{\gamma_{R}}D_{R}^{*},$
-$D_{A}^{*}$ $\xrightarrow{\alpha_{a}^{\prime}}M_{A},$ $D_{R}^{*}$ $\xrightarrow{\alpha_{r}^{\prime}}M_{R},$
-$M_{A}$ $\xrightarrow{\delta_{ma}}\phi,$ $M_{R}$ $\xrightarrow{\delta_{mr}}\phi,$
+Figure 6: Time series trajectories for three key species in the Vilar oscillator system: Complex (C), Activator (A), and Repressor (R).
+The darker line shows a reference trajectory obtained using the true parameter values, while the four lighter lines represent additional stochastic realizations under identical conditions.
+This visualization highlights the intrinsic variability of the system while preserving its consistent oscillatory behavior.
+The network includes the following reactions:
+$D_{A},A$ $\xrightarrow{\gamma_{A}}D_{A}^{*},$ $D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},$
+$D_{A}^{*}$ $\xrightarrow{\theta_{A}}D_{A},\,A,$ $D_{A}^{*}$ $\xrightarrow{\alpha_{A}^{*}}D_{A}^{*},\,M_{A},$
+$D_{A}$ $\xrightarrow{\alpha_{A}}D_{A},\,M_{A},$
+$D_{R},A$ $\xrightarrow{\gamma_{R}}D_{R}^{*},$ $D_{R}^{*}$ $\xrightarrow{\theta_{R}}D_{R},$
+$D_{R}^{*}$ $\xrightarrow{\theta_{A}}D_{R}^{*},\,A,$ $D_{R}^{*}$ $\xrightarrow{\alpha_{R}^{*}}D_{R}^{*},\,M_{R},$
+$D_{R}$ $\xrightarrow{\alpha_{R}}D_{R},\,M_{R},$
 $M_{A}$ $\xrightarrow{\beta_{A}}M_{A}+A,$ $M_{R}$ $\xrightarrow{\beta_{R}}M_{R}+R,$
+$M_{A}$ $\xrightarrow{\delta_{M_{A}}}\phi,$ $M_{R}$ $\xrightarrow{\delta_{M_{R}}}\phi,$
 $A,R$ $\xrightarrow{\gamma_{C}}C,$ $C$ $\xrightarrow{\delta_{A}}R,$
@@
   • True parameters:
-    $\alpha_{a}$ $=50,\quad\alpha_{a}^{\prime}=500,\quad\alpha_{r}=0.01,\quad\alpha_{r}^{\prime}=50,$
-    $\beta_{a}$ $=50,\quad\beta_{r}=5,\quad\delta_{ma}=10,\quad\delta_{mr}=0.5,$
-    $\delta_{a}$ $=1,\quad\delta_{r}=0.2,\quad\gamma_{a}=1,\quad\gamma_{r}=1,$
-    $\gamma_{c}$ $=2,\quad\theta_{a}=50,\quad\theta_{r}=100.$
+[…]
+[…]
+[…]
+[…]
   • Initial Species Values:
     $D_{A}=1,\;D_{A}^{*}=0,\;M_{A}=0,\;D_{R}=1,\;D_{R}^{*}=0,\;M_{R}=0,\;C=10,\;A=10,\;R=10.$
-[…]
-Figure 7: Time series trajectories for three key species in the Vilar oscillator system: Complex (C), Activator (A), and Repressor (R).
-The darker line shows a reference trajectory obtained using the true parameter values, while the four lighter lines represent additional stochastic realizations under identical conditions.
-This visualization highlights the intrinsic variability of the system while preserving its consistent oscillatory behavior.
-Appendix B Hyperparameters
-In this section, we describe the hyperparameters and simulation environments employed across all methods.
-Training data for the SBI benchmark problems were generated using the sbibm [Lueckmann et al., 2021], except for the LV and SIR models, which were generated with the BayesFlow framework [Radev et al., 2023b].
-For the HH model, we adopt the implementation provided by [Gloeckler et al., 2024], and for the Genetic Oscillator model, we follow the approach of [Akesson et al., 2022].
-In this model, the simulated data is high dimensional, with a shape of $(N\times 3\times 200)$ representing three species observed over 200 timesteps.
-We compress this data using an autoencoder whose latent space is set to match the number of simulator parameters.
-For Simformer, we use hyperparameters published in their original works [Gloeckler et al., 2024].
-For GATSBI, task-specific hyperparameters were taken from [Ramesh et al., 2022] and adjusted as needed based on each problem’s structure and computational demands.
-For NPE, we follow [Lueckmann et al., 2021], which evaluated NPE across ten benchmark problems.
-As reported in Appendix H of sbibm paper, Neural Spline Flows (NSF) generally outperform Masked Autoregressive Flows (MAF), originally recommended for NPE.
-Consequently, we use the NSF implementation from the sbi package [Tejero-Cantero et al., 2020] as the density estimator in all experiments.
-The implementation of Automatic Posterior Transformation (APT, or SNPE-C) is also taken from the sbi package; as a sequential method, the simulation budget per round is computed by dividing the total budget by the number of rounds.
-For time series problems, a summary network is recommended for NPE/APT [Zeghal et al., 2022].
-Accordingly, we modify and optimize the standard CNNEmbedding from the sbi package to convert the data into summary statistics with dimensions matching the parameters (4 for the LV test problem and 2 for the SIR task).
-For the ConDiSim model, we optimized the hyperparameters for each of the ten benchmark problems from the sbibm library, as detailed in Table 3.
-Common hyperparameters were applied uniformly across all tasks to ensure training consistency and stability.
-We used the AdamW optimizer with a learning rate of $0.001$ and a weight decay of $0.0001$, along with a ReduceLROnPlateau scheduler to decrease the learning rate by a factor of $0.5$ if validation loss did not improve over five consecutive epochs.
-The noise schedule followed a cosine/quadratic scheme, with $\beta_{\text{min}}$ as $0.0001$ and $\beta_{\text{max}}$ as $0.02$, to regulate progressive noise addition during diffusion.
-Data standardization was conducted using StandardScaler, and an $80/20$ split was used for training and validation.
-Early stopping with a patience of $20$ epochs was applied to prevent overfitting.
-SiLU (Sigmoid Linear Unit) activations were used across all layers to enhance pattern capture and improve stability.
-This consistent training strategy enabled a systematic comparison of model performance across various batch sizes, diffusion steps, and hidden unit configurations.
-Table 3: Hyperparameters for ConDiSim Model
-Benchmark Batch Size Diffusion Steps Hidden Units SLCP 32 200 Linear, (512, 512, 512) SLCP with Distractors 32 1000 Linear, (512, 512, 512) Two Moons 32 50 Linear, (256, 256, 256) Gaussian Mixture 32 50 Linear, (256, 256, 256) Gaussian Linear Uniform 32 50 Linear, (256, 256, 256) SIR 32 50 Linear, (256, 256, 256) Bernoulli GLM 128 500 Linear, (256, 256, 256) Bernoulli GLM Raw 128 500 Linear, (512, 512, 512) Lotka-Volterra 128 200 Conv1D (32, 64, 128), Linear, (128, 128) HH Model 32 300 Linear, (256, 256, 256) Genetic Oscillator 64 400 Linear, (256, 256, 256, 256)
-Appendix C Additional Results
-We present plots of estimated posterior distributions for the remaining benchmark test problems below:
-C.1 Posterior Evolution for Two Moons
+  • Simulation Time: The system is simulated over $200$ time steps uniformly distributed across $200\,\text{ms}$.
+Appendix C Hyperparameters
+In this section, we describe the hyperparameters and simulation environments used across all methods.
+Training data for the SBI benchmark problems were generated using sbibm [Lueckmann et al., 2021], except for the Lotka–Volterra (LV) and SIR models, which were simulated with BayesFlow [Radev et al., 2023b].
+For SimFormer, we adopt the hyperparameters reported in Gloeckler et al.
+[2024] for both the general SBI setting and the Hodgkin–Huxley (HH) model.
+For the Genetic Oscillator model, we follow the simulation and preprocessing setup of Akesson et al.
+[2022].
+Specifically, SimFormer is trained on a simulation budget of 30,000 using an 8-layer Transformer with 8 attention heads, 64-dimensional head projections, a 512-unit MLP, and dropout rate 0.1; it is optimized with Adam (learning rate $10^{-3}$) for 100,000 iterations with batch size 128, and posterior samples are generated via Euler–Maruyama integration over 1,000 diffusion steps.
+For GATSBI, we start from the task-specific hyperparameters in Ramesh et al.
+[2022] and adjust them as needed based on problem structure and computational constraints.
+All experiments were run on a system with dual 16-core Intel Xeon Gold 6226R CPUs, 576 GB RAM, and an NVIDIA Tesla T4 GPU (16 GB).
+For NPE, we adopt the setup of Lueckmann et al.
+[2021], who evaluate NPE across ten sbibm benchmark tasks.
+Appendix H of their work shows that Neural Spline Flows (NSF) consistently outperform Masked Autoregressive Flows (MAF), which were originally proposed as the density estimator for NPE.
+We therefore use the NSF implementation from the sbi package [Tejero-Cantero et al., 2020] in all experiments.
+We also use sbi’s implementation of Automatic Posterior Transformation (APT, or SNPE-C).
+As a sequential method, APT allocates the total simulation budget uniformly across rounds.
+For ConDiSim, we tuned hyperparameters individually for each task in the sbibm benchmark suite [Lueckmann et al., 2021] while keeping the training pipeline fixed.
+We optimized with AdamW [Loshchilov and Hutter, 2019] using a YOLOX-inspired learning rate schedule: a quadratic warmup over the first $10\%$ of steps from $0.2\eta$ to the base learning rate $\eta$, followed by cosine annealing down to $10^{-6}$, and a constant phase at $10^{-6}$ for the final $5\%$ of steps; weight decay was set to $\lambda=10^{-4}$.
+We use scaled linear and scaled quadratic noise schedule [Ho et al., 2020] with $\beta_{\text{start}}=10^{-4}s$ and $\beta_{\text{end}}=0.02s$ with $s=1000/T$ to maintain consistent noise levels across varying diffusion steps $T$.
+Inputs were standardized to zero mean and unit variance, and a $70/30$ train–validation split was employed to handle the high variance of posterior estimation.
+Training stability was ensured via gradient clipping (max norm $5.0$) and SNR-based loss weighting [Hang et al., 2023] with $\gamma=5.0$, where timestep weights $w_{t}=\min(\mathrm{SNR}_{t},\gamma)/\mathrm{SNR}_{t}$ and $\mathrm{SNR}_{t}=\bar{\alpha}_{t}/(1-\bar{\alpha}_{t})$ emphasize learning at noisier timesteps.
+Early stopping (patience $30$ after epoch $50$) and SiLU activations throughout the network were also used.
+Task-specific hyperparameters are provided in Table 3.
+Table 3: ConDiSim hyperparameters.
+Benchmark # Diffusion Blocks Hidden Dim Scheduler # Timesteps Batch Size Learning Rate SLCP 6 128 linear 200 32 1e-3 SLCP with Distractors 6 128 quadratic 1000 50 1e-3 Two Moons 4 64 cosine 160 32 1e-3 Gaussian Mixture 4 64 cosine 160 32 1e-3 Gaussian Linear 6 64 linear 100 50 2e-4 Gaussian Linear Uniform 6 64 cosine 200 50 2e-4 SIR 4 64 linear 100 32 1e-4 Bernoulli GLM 6 128 linear 200 32 1e-3 Bernoulli GLM Raw 6 128 linear 200 32 1e-3 Lotka-Volterra 6 128 cosine 200 50 2e-4 HH Model 6 128 quadratic 1000 50 1e-3 Genetic Oscillator 6 128 cosine 300 64 1e-4
+Appendix D Additional Results
+D.1 Posterior Evolution for Two Moons
 The evolution of the approximated posterior for the Two Moons test problem is shown in Fig. 7.
 Figure 7: Evolution of the posterior distribution for the Two Moons task, showing the transition from a broad, noisy distribution to a well-defined bi-modal structure as $T\to 0$.
-(a) Hodgkin–Huxley: Approximated Posterior
-(b) Posterior Predictive Voltage Traces
-Figure 9: Simformer:Posterior Distributions for Hodgkin–Huxley model
-C.2 Performance of Simformer on HH Model
-Unlike the standard SBI benchmarks, where true posteriors are known for a known set of true parameters, the Hodgkin-Huxley (HH) model lacks an analytical posterior, preventing direct comparison via metrics like C2ST or MMD.
-For fair evaluation, we first benchmark CondiSim on 10 standard SBI tasks (refer Section 4.1) before demonstrating its real-world applicability to the HH model.
-Given a fixed parameter configuration $\theta^{*}=(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606)$, we infer posteriors using both CondiSim and Simformer (trained with 30k simulations).
-Posterior samples and posterior predictive checks for Simformer (Figure 9) and ConDiSim (Figure 4) confirm that both methods produce plausible voltage traces, with ConDiSim yielding tighter posterior support on several parameters.
-Appendix D Posterior Plots for SBI benchmarking problems
-In this section, we summarize the unique challenges posed by remaining SBI benchmark problem and highlight the performance of our model.
+D.2 Comparison with Simformer on HH and Genetic Oscillator Models
+Unlike standard SBI benchmarks with analytical posteriors, the Hodgkin–Huxley (HH) and Genetic Oscillator models lack closed-form solutions, preventing direct evaluation via C2ST or MMD.
+Having benchmarked ConDiSim on ten standard SBI tasks (Section 4.1), we now compare it against Simformer, the closest competing method, on these real-world problems.
+Both methods are trained with 30k simulations.
+For the HH model with fixed parameters $\bm{\theta^{*}}=(1.437,72.177,24.209,0.154,66.87,-81.435,-67.606)$, posterior samples and predictive checks (Figures 84) show both methods recover plausible voltage traces, with ConDiSim yielding tighter posterior support for several parameters.
+For the genetic oscillator, posterior distributions for Simformer and ConDiSim are shown in Figures 9 and 5(a), respectively.
+Both produce reasonable posteriors, but ConDiSim achieves sharper concentration around true parameters for several rates (e.g., $\alpha_{r},\beta_{r},\gamma_{c},\theta_{a}$), while Simformer yields broader supports.
+This demonstrates ConDiSim’s more precise uncertainty quantification without sacrificing coverage, particularly effective for complex oscillatory dynamics.
+(a) Hodgkin–Huxley: Approximated Posterior.
+(b) Posterior Predictive Voltage Traces.
+Figure 8: Simformer: Posterior Distributions for Hodgkin–Huxley model.
+Figure 9: Simformer: Posterior Distributions for the Genetic Oscillator test problem.
+Appendix E Posterior Plots for SBI benchmarking problems
+In this section, we summarize the unique challenges posed by the remaining SBI benchmark problem and highlight the performance of our model.
 Detailed posterior distributions are provided in the corresponding figures for visual reference.
@@
 The Gaussian Linear Uniform variant complicates posterior inference with a uniform prior.
-Figure 11(a) and Figure 11(b) shows that our model effectively handles these differences.
+Figure 11(a) and Figure 11(b) show that our model effectively handles these differences.
 In Bernoulli GLM and Bernoulli GLM Raw, the challenge involves estimating posteriors from Bernoulli observations with Gaussian priors.
-The "Raw" variant introduces complexity by using raw data instead of sufficient statistics.
+The “Raw” variant introduces complexity by using raw data instead of sufficient statistics.
 Figures 12(a) and 12(b) indicate accurate posterior modeling for both cases.
 The SIR model aims to infer contact and recovery rates from noisy infection data over time.
-Figure 10(b) illustrates that our model robustly estimates the parameters despite the high variability in data.
+Figure 10(b) illustrates that our model robustly estimates the parameters despite the high variability in the data.
 Lastly, the Lotka-Volterra model is challenging due to its nonlinear dynamics and potential chaotic behavior.
@@
 These experiments demonstrate that ConDiSim effectively infers posterior distributions across all benchmark problems with high accuracy.
-(a) GMM
-(b) SIR Model
+(a) GMM.
+(b) SIR Model.
 Figure 10: Posterior Plots: Comparison of generated vs. reference posterior distributions for the GMM and SIR models.
-(a) Gaussian Linear Model
-(b) Gaussian Linear Uniform Model
-Figure 11: Posterior Distributions for Gaussian Linear and Gaussian Linear Uniform Models: Generated vs. Reference Posterior
-(a) Bernoulli GLM
-(b) Bernoulli GLM Raw
-Figure 12: Posterior Distributions for Bernoulli GLM and Bernoulli GLM Raw: Generated vs. Reference Posterior
-Figure 13: Lotka–Volterra Posterior: Visualization of the posterior for the Lotka–Volterra model. The true posterior mean is indicated by a vertical line (due to low variance).
-Appendix E Simulation Based Calibration
-Simulation-Based Calibration (SBC) is an essential metric for generative models in Simulation-Based Inference (SBI), as it verifies alignment between inferred posterior distributions and true parameter values, assisting in detecting potential biases in model-generated samples.
+(a) Gaussian Linear Model.
+(b) Gaussian Linear Uniform Model.
+Figure 11: Posterior Distributions for Gaussian Linear and Gaussian Linear Uniform Models: Generated vs. Reference Posterior.
+(a) Bernoulli GLM.
+(b) Bernoulli GLM Raw.
+Figure 12: Posterior Distributions for Bernoulli GLM and Bernoulli GLM Raw: Generated vs. Reference Posterior.
+Figure 13: Lotka Volterra Posterior: Visualization of the posterior for the Lotka–Volterra model. The true posterior mean is indicated by a vertical line (due to low variance).
+Appendix F Simulation Based Calibration
+Simulation-Based Calibration (SBC) is an important metric for generative models in SBI, as it verifies alignment between inferred posterior distributions and true parameter values, assisting in detecting potential biases in model-generated samples.
 For SBC, we generate $M=500$ prior samples, simulate corresponding observations, and draw $L=250$ posterior samples for each observation using the learned ConDiSim model.
@@
 The SBC results for SBI benchmarking problems are presented below:
-(a) SLCP
-(b) SLCP Distractors
-(c) Gaussian Linear
-(d) Gaussian Linear Uniform
-(e) Bernoulli GLM
-(f) Bernoulli GLM Raw
-(g) SIR
-(h) Gaussian Mixture
-(i) Lotka Volterra
+(a) SLCP.
+(b) SLCP Distractors.
+(c) Gaussian Linear.
+(d) Gaussian Linear Uniform.
+(e) Bernoulli GLM.
+(f) Bernoulli GLM Raw.
+(g) SIR.
+(h) Gaussian Mixture.
+(i) Lotka Volterra.
 Figure 15: ECDF results for the 30,000 simulation budget and the first run for each problem.
-Appendix F SBI Metrics
-F.1 Limitations of the C2ST Metric
+Appendix G SBI Metrics
+G.1 Limitations of the C2ST Metric
 The Classifier Two-Sample Test (C2ST) evaluates the similarity between a generated posterior $P$ and the true posterior $Q$ by training a classifier to differentiate their samples.
 A score near 0.5 indicates that the two distributions are similar, while higher values indicate discrepancies.
-Although widely used in SBI literature for its straightforward interpretability—providing a clear score between 0.5 and 1—the metric can be overly sensitive, as a single poorly estimated dimension may distort the results.
-Furthermore, its reliability depends significantly on the design and training of the classifier.
-F.2 MMD
-To overcome these limitations, we also consider the Maximum Mean Discrepancy (MMD), which measures the distance between the mean embeddings of $P$ and $Q$ in a reproducing kernel Hilbert space (RKHS).
-MMD is defined as,
+Although widely used in SBI literature for its straightforward interpretability providing a clear score between 0.5 and 1, the metric can be overly sensitive, as a single poorly estimated dimension may distort the results.
+Furthermore, its reliability depends significantly on the design and training of the classifier [Friedman, 2003, Lopez-Paz and Oquab, 2016].
+However, C2ST remains the most widely used metric in SBI, as it is both interpretable and well established for evaluating benchmark problems.
+For completeness, we also report MMD results in the following section, although MMD is kernel dependent, less commonly used, and often difficult to interpret solely from numerical values.
+G.2 MMD
+The Maximum Mean Discrepancy (MMD) measures the distance between the mean embeddings of $P$ and $Q$ in a reproducing kernel Hilbert space (RKHS). MMD is defined as,
 $\mathrm{MMD}^{2}(P,Q;\mathcal{H})=\left\|\mathbb{E}_{x\sim P}[\phi(x)]-\mathbb{E}_{y\sim Q}[\phi(y)]\right\|_{\mathcal{H}}^{2},$ (38)
@@
 A low MMD value indicates that the generated posterior closely approximates the true posterior.
-Unlike C2ST, MMD provides a robust, holistic measure of similarity between distributions without being overly influenced by discrepancies in a single dimension.
+MMD therefore provides a holistic measure of similarity between distributions without being overly influenced by discrepancies in a single dimension.
 A low MMD value indicates that the distributions $P$ and $Q$ are similar, while a larger value suggests significant differences.
 Table 4: MMD metric for different methods across simulation budgets (mean $\pm$ std. dev.).
-(a) Two Moons, Bernoulli GLM, and SIR
-Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 FMPE 0.0133 $\pm$ 0.0197 0.0039 $\pm$ 0.0060 0.019 $\pm$ 0.0014 0.1076 $\pm$ 0.0283 0.0826 $\pm$ 0.0138 0.0746 $\pm$ 0.0247 0.2058 $\pm$ 0.0024 0.3618 $\pm$ 0.0048 0.1819 $\pm$ 0.0044 ConDiSim 0.0105 $\pm$0.0041 0.0104 $\pm$0.0048 0.0074 $\pm$0.0002 0.3157 $\pm$0.1771 0.1471 $\pm$0.0428 0.1981 $\pm$0.1348 0.7560 $\pm$0.1803 0.5754 $\pm$0.2392 0.6909 $\pm$0.1521
-(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra
-Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 FMPE 0.1598 $\pm$ 0.0501 0.1403 $\pm$ 0.0471 0.0930 $\pm$ 0.0561 0.0859 $\pm$ 0.0167 0.0831 $\pm$ 0.0181 0.0781 $\pm$ 0.0102 1.0537 $\pm$ 0.6277 1.0858 $\pm$ 0.5639 1.1696 $\pm$ 0.4336 ConDiSim 1.0688 $\pm$0.3057 0.6939 $\pm$0.3204 0.4105 $\pm$0.3082 0.1448 $\pm$0.0368 0.1435 $\pm$0.0822 0.1658 $\pm$0.1407 3.4480 $\pm$0.8115 3.4678 $\pm$0.3850 3.7210 $\pm$0.5691
-(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP
-Method Gaussian Mixture Gaussian Linear Gaussian Linear Uniform SLCP Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 GATSBI 0.2242 $\pm$0.0296 0.1149 $\pm$0.0549 0.1733 $\pm$0.1082 0.4999 $\pm$0.1595 0.6476 $\pm$0.3050 0.7182 $\pm$0.2839 0.9044 $\pm$0.2353 0.7712 $\pm$0.1483 1.1197 $\pm$0.4795 0.7197 $\pm$0.5598 0.3375 $\pm$0.1473 0.7948 $\pm$0.5992 NPE 0.0604 $\pm$0.0260 0.0418 $\pm$0.0211 0.0554 $\pm$0.0347 0.0105 $\pm$0.0036 0.0074 $\pm$0.0018 0.0082 $\pm$0.0011 0.0258 $\pm$0.0098 0.0117 $\pm$0.0041 0.0121 $\pm$0.0051 0.4021 $\pm$0.0551 0.1632 $\pm$0.0684 0.2333 $\pm$0.0935 APT 0.0137 $\pm$0.0031 0.0108 $\pm$0.0088 0.0110 $\pm$0.0034 0.0075 $\pm$0.0013 0.0070 $\pm$0.0018 0.0080 $\pm$0.0008 0.0166 $\pm$0.0087 0.0138 $\pm$0.0053 0.0113 $\pm$0.0039 0.1405 $\pm$0.0347 0.0487 $\pm$0.0296 0.0287 $\pm$0.0166 Simformer 0.0009 $\pm$0.0007 0.0009 $\pm$0.0006 0.0007 $\pm$0.0010 0.0092 $\pm$0.0116 0.0037 $\pm$0.0035 0.0037 $\pm$0.0028 0.0014 $\pm$0.0006 0.0015 $\pm$0.0011 0.0008 $\pm$0.0003 0.0323 $\pm$0.0203 0.0025 $\pm$0.0034 0.0021 $\pm$0.0030 FMPE 0.1893 $\pm$ 0.0985 0.0208 $\pm$ 0.0078 0.0013 $\pm$ 0.0011 0.0511 $\pm$ 0.0053 $0$$\pm$ 0.0041 0.0497 $\pm$ 0.0034 0.1425 $\pm$ 0.0219 0.1261 $\pm$ 0.0206 0.1236 $\pm$ 0.0247 0.1573 $\pm$ 0.0543 0.1489 $\pm$ 0.0528 0.1489 $\pm$ 0.0528 ConDiSim 0.1604 $\pm$0.0599 0.0720 $\pm$0.0124 0.0974 $\pm$0.0505 0.0471 $\pm$0.0126 0.0387 $\pm$0.0129 0.0290 $\pm$0.0038 0.0335$\pm$0.0139 0.0273 $\pm$0.0079 0.0285 $\pm$0.0098 0.2760 $\pm$0.0493 0.1954 $\pm$0.1029 0.1572 $\pm$0.0726
+(a) Two Moons, Bernoulli GLM, and SIR.
+Method Two Moons Bernoulli GLM SIR Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 0.0254 $\pm$0.0025 0.0241 $\pm$0.0099 0.0125 $\pm$0.0045 1.0371 $\pm$0.1830 0.9379 $\pm$0.1001 0.8675 $\pm$0.1945 0.6611 $\pm$0.2236 1.0312 $\pm$0.1255 1.1903 $\pm$0.1007 APT 0.0073 $\pm$0.0030 0.0069 $\pm$0.0030 0.0037 $\pm$0.0012 0.7383 $\pm$0.0853 0.4944 $\pm$0.0663 0.5808 $\pm$0.0539 1.1003 $\pm$0.0712 1.1093 $\pm$0.0311 1.1291 $\pm$0.0481 GATSBI 0.1213 $\pm$0.0650 0.0582 $\pm$0.0122 0.1432 $\pm$0.1161 3.6381 $\pm$0.5428 4.8299 $\pm$0.6825 3.9965 $\pm$1.8510 2.9470 $\pm$0.4891 2.7502 $\pm$0.7268 2.1466 $\pm$0.5625 Simformer 0.0005 $\pm$0.0003 0.0001 $\pm$0.0001 0.0003 $\pm$0.0004 0.0191 $\pm$0.0098 0.0156 $\pm$0.0087 0.0116 $\pm$0.0061 0.0018 $\pm$0.0028 0.0016 $\pm$0.0042 0.0016 $\pm$0.0085 FMPE 0.0133 $\pm$ 0.0197 0.0039 $\pm$ 0.0060 0.019 $\pm$ 0.0014 0.1076 $\pm$ 0.0283 0.0826 $\pm$ 0.0138 0.0746 $\pm$ 0.0247 0.2058 $\pm$ 0.0024 0.3618 $\pm$ 0.0048 0.1819 $\pm$ 0.0044 ConDiSim 0.0172 $\pm$0.0226 0.0091 $\pm$0.0099 0.0052 $\pm$0.0018 0.0393 $\pm$0.0095 0.0525 $\pm$0.0547 0.0190 $\pm$0.0134 0.2413 $\pm$0.0681 0.3853 $\pm$0.1195 0.4860 $\pm$0.1198
+(b) SLCP Distractors, Bernoulli GLM Raw, and Lotka Volterra.
+Method SLCP Distractors Bernoulli GLM Raw Lotka Volterra Budget 10,000 20,000 30,000 10,000 20,000 30,000 10,000 20,000 30,000 NPE 1.2167 $\pm$0.2397 0.7834 $\pm$0.3453 0.6428 $\pm$0.1779 1.1227 $\pm$0.0987 1.1389 $\pm$0.1878 1.1448 $\pm$0.1287 2.6357 $\pm$0.3468 3.8635 $\pm$0.2770 4.3310 $\pm$0.0709 APT 1.8380 $\pm$0.4448 0.8278 $\pm$0.2373 0.4422 $\pm$0.1428 0.7233 $\pm$0.2131 0.6968 $\pm$0.0826 0.5437 $\pm$0.1333 3.1921 $\pm$0.8198 3.2722 $\pm$0.8169 4.4739 $\pm$0.9642 GATSBI 5.1993 $\pm$1.2320 7.2478 $\pm$1.2260 3.0631 $\pm$0.3244 3.4369 $\pm$0.8517 4.3479 $\pm$0.5433 3.5950 $\pm$0.8613 7.2684 $\pm$0.5733 5.5183 $\pm$1.7126 4.8858 $\pm$1.2146 Simformer 0.0540 $\pm$0.0173 0.0541 $\pm$0.0161 0.0552 $\pm$0.0195 0.2056 $\pm$0.0486 0.0926 $\pm$0.0493 0.0731 $\pm$0.0238 0.9696 $\pm$0.4434 0.8342 $\pm$0.4742 0.9634 $\pm$0.2588 FMPE 0.1598 $\pm$ 0.0501 0.1403 $\pm$ 0.0471 0.0930 $\pm$ 0.0561 0.0859 $\pm$ 0.0167 0.0831 $\pm$ 0.0181 0.0781 $\pm$ 0.0102 1.0537 $\pm$ 0.6277 1.0858 $\pm$ 0.5639 1.1696 $\pm$ 0.4336 ConDiSim 0.1378 $\pm$0.0404 0.0580 $\pm$0.0205 0.0275 $\pm$0.0182 0.0968 $\pm$0.1012 0.0359 $\pm$0.0212 0.0305 $\pm$0.0279 2.3226 $\pm$0.1486 2.1965 $\pm$0.3856 2.1180 $\pm$0.2584
+(c) Gaussian Mixture, Gaussian Linear, Gaussian Linear Uniform, and SLCP.
+[…]